### v1

# Written by AAAI Press Staff1

Abstract
Graph partitioning aims to divide a graph into $k$ disjoint subsets while optimizing a specific partitioning objective. The majority of formulations related to graph partitioning exhibit NP-hardness due to their combinatorial nature. As a result, conventional approximation algorithms rely on heuristic methods, sometimes with approximation guarantees and sometimes without. Unfortunately, traditional approaches are tailored for specific partitioning objectives and do not generalize well across other known partitioning objectives from the literature. To overcome this limitation, and learn heuristics from the data directly, neural approaches have emerged, demonstrating promising outcomes. In this study, we extend this line of work through a novel framework, NeuroCUT. NeuroCUT introduces two key innovations over prevailing methodologies. First, it is inductive to both graph topology and the partition count, which is provided at query time. Second, by leveraging a reinforcement learning based framework over node representations derived from a graph neural network, NeuroCUT can accommodate any optimization objective, even those encompassing non-differentiable functions. Through empirical evaluation, we demonstrate that NeuroCUT excels in identifying high-quality partitions, showcases strong generalization across a wide spectrum of partitioning objectives, and exhibits resilience to topological modifications.
Abstract
1 Introduction and Related Work
Graph partitioning is a fundamental problem in network analysis with numerous real-world applications in various domains such as system design in online social networks (pujol2009divide), dynamic ride-sharing in transportation systems (tafreshian2020trip), VLSI design (kahng2011vlsi), and preventing cascading failure in power grids (li2005strategic). The goal of graph partitioning is to divide a given graph into disjoint subsets where nodes within each subset exhibit strong internal connections while having limited connections with nodes in other subsets. Generally speaking, the aim is to find somewhat balanced partitions while minimizing the number of edges across partitions.
Several graph partitioning formulations have been studied in the literature, mostly in the form of discrete optimization (karypis1997multilevel; karypis1998multilevelk; karypis1999multilevel; andersen2006local; chung2007four). The majority of the formulations are NP-hard and thus the proposed solutions are either heuristics or algorithms with approximate solutions (karypis1997multilevel).
In recent years there has been a growing interest in using machine learning techniques to design heuristics for NP-Hard combinatorial problems on graphs. This line of work is motivated from the observation that traditional algorithmic approaches are limited to using the same heuristic irrespective of the underlying data distribution. Neural approaches, instead, focus on learning the heuristic as a function of the data without deriving explicit algorithms (duong2023deep; jung2022learning; nazi2019gap). Significant advances have been made in learning to solve a variety of combinatorial optimization problems such as Set Cover, Travelling Salesman Problem, Influence Maximization etc. (khalil2017learning; gcomb; kool2018attention; joshi2019efficient). The neural architectural designs in these methods are generally suitable for a specific set of CO problems (mazyavkina2021reinforcement) of similar nature. For instance, AM  (kool2018attention) and NeuroLKH (xin2021neurolkh) are targeted towards routing problems. GCOMB (gcomb) and LeNSE (ireland2022lense) aim to solve problems that are submodular in nature. Owing to these assumptions, the architectural designs in these works are not suitable for our problem of $k$-way inductive graph partitioning with versatile partitioning objectives.
The most relevant method to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (nazi2019gap). It proposes a differentiable loss function for partitioning based on a continuous relaxation of the normalized cut formulation. Deep-MinCut being an unsupervised approach learns both node embeddings and the community structures simultaneously where the objective is to minimize the mincut loss (duong2023deep). Another method solves the multicut problem where the number of partitions is not an input to the problem (jung2022learning). The idea is to construct a reformulation of the multicut ILP constraints to a polynomial program as a loss function. Finally,  (gatti2022graph) solves the normalized cut problem only for the case where the number of partitions is exactly two.
Nevertheless, existing neural approaches that tackle graph partitioning problem suffer three notable drawbacks, which we address.
  • Assumption of differential objective function: Existing neural approaches assume the objective function to be differentiable. As we will illustrate in § 2, the assumption does not always hold in the real-world.
  • Inductivity to partition count: The number of partitions required to segment a graph is an input parameter. Hence, it is important for a learned model to generalize to any partition count without retraining. Existing neural approaches are non-inductive to the number of partitions, i.e they can only infer on number of partitions on which they are trained. Additionally, it’s worth noting that the optimal number of partitions is often unknown beforehand. Hence, it is a common practice to experiment with different partition counts and evaluate their impact on the partitioning objective. Typically, the optimal partition count corresponds to the point where the second-order derivative of the objective function approaches saturation. Therefore, inductivity to the number of partitions becomes an essential requirement for conducting such experiments in a scalable manner.
  • Generalizability to objective functions: Multiple objective functions have been studied in the partitioning literature. The optimal objective function hinges upon the subsequent application in question. Consequently, it holds significance for an algorithmic framework to possess the capacity to encompass a diverse array of objective functions. Unfortunately, prevailing neural frameworks are founded upon presumptions that yield efficacy exclusively within the confines of their designated objective function.
1.1 Contributions:
In this paper, we circumvent the above-mentioned limitations through the following key contributions.
  • Versatile objectives: We develop an auto-regressive, graph reinforcement learning framework, called NeuroCUT, to solve the graph partitioning problem. Diverging from conventional algorithms, NeuroCUT can handle different partitioning objectives. Moreover, unlike other neural methods, NeuroCUT can accommodate diverse partitioning objectives, without the necessity for differentiability.
  • Inductivity to topology and number of partitions: The parameter space of NeuroCUT is independent of the graph size or the partition count. This innovative architecture endows NeuroCUT with the ability to generalize effectively to previously unseen graphs of varying scales and adapt to arbitrary partition count specified during inference.
  • Empirical Assessment: We perform comprehensive experiments employing real-world datasets, evaluating NeuroCUT across four distinct graph partitioning objectives. Our empirical investigation substantiates the efficacy of NeuroCUT in partitioning tasks, showcasing its robustness across a spectrum of objective functions, and demonstrating its capacity to generalize effectively to partition sizes and topological configurations that it has not encountered during training.
2 Problem Formulation
In this section, we introduce the concepts central to our work and formulate the problem. All the notations used in this work are outlined in Tab. 1 in appendix.
Definition 1 (Graph)
We denote a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{G}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges.
Definition 2 (Cut)
A cut $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is a partition of $\mathcal{V}$ into two subsets $\mathcal{S}$ and $\mathcal{T}$. The cut-set of $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is the set $(u,v)\in\mathcal{E}|u\in\mathcal{S},v\in\mathcal{T}$ of edges that have one endpoint in $\mathcal{S}$ and the other endpoint in $\mathcal{T}$.
Definition 3 (Graph Partitioning)
Given graph $\mathcal{G}$, we aim to partition $\mathcal{G}$ into $k$ disjoint sets $\mathcal{P}=\{P_{1},P_{2},\cdots,P_{k}\}$ such that the union of the nodes in those sets is equal to $\mathcal{V}$ i.e $\cup_{i=1}^{k}P_{i}=\mathcal{V}$ and each node belongs to exactly one partition.
Partitioning Objective: We aim to minimize/maximize a partitioning objective of the form ${Obj}(\mathcal{G},\mathcal{P})$. A wide variety of objectives for graph partitioning have been proposed in the literature. Without loss of generality, we consider the following four objectives. These objectives are chosen due to being well studied in the literature, while also being diverse from each other.  [^1] Our framework is not restricted to these objectives.
  1. 1.
    $k$-MinCut (saran1995finding): Partition a graph into $k$ partitions such that the total number of edges across partitions is minimized.
[…]
    Here $P_{l}$ refers to the set of elements in $l^{th}$ partition of $\mathcal{P}$ as described in Def.  3 and $\overline{P_{l}}$ refers to set of elements not in $P_{l}$.
  2. 2.
    Normalized Mincut (shi2000normalized): The $k$-minimum cut criteria favors cutting small sets of isolated nodes in the graph. To avoid this unnatural bias for partitioning out small sets of points, normalised cut computes the cut cost as a fraction of the total edge connections to all the nodes in the graph.
    $\operatorname{Ncut}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}$ (2)
    Here, $\operatorname{vol}\left(P_{l},\mathcal{V}\right):=\sum_{v_{i}\in P_{l},v_{j}\in\mathcal{V}}e(v_{i},v_{j})$.
  3. 3.
    Balanced Cuts (nazi2019gap): Balanced cut favours partitions of equal sizes so an extra term that indicates the squared distance from equal sized partition is added to normalised cuts.
    $\operatorname{Balanced-Cuts}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}+\frac{(|P_{l}|-|\mathcal{V}|/k)^{2}}{|\mathcal{V}|^{2}}$ (3)
    Here, $\operatorname{vol}\left(P_{l},\mathcal{V}\right):=\sum_{v_{i}\in P_{l},v_{j}\in\mathcal{V}}e(v_{i},v_{j})$
  4. 4.
    (Weighted) Sparsest Cuts (chawla2006hardness): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition. We generalize it to $k$-way sparsest cuts by summing up the value for all the partitions. The intuition behind sparsest cuts is that any partition should neither be very large nor very small.
    $\phi(S,\bar{S})$ $=\frac{\operatorname{cut}(S,\bar{S})}{\min(|S|,|\bar{S}|)}$ (4)
    $\operatorname{Sparsest-Cuts}(\mathcal{\mathcal{P}})$ $=\sum_{l=1}^{|\mathcal{P}|}\phi(P_{l},\overline{P_{l}})$ (5)
Problem 1 (Learning to Partition Graph )
Given a graph $\mathcal{G}$ and the number of partitions $k$, the goal is to find a partitioning $\mathcal{P}$ of the graph $\mathcal{G}$ that optimizes a target objective function ${Obj}(\mathcal{G},\mathcal{P})$. Towards this end, we aim to learn a policy $\pi$ that assigns each node $v\in\mathcal{V}$ to a partition in $\mathcal{P}$.
In addition to our primary goal of finding a partitioning that optimizes a certain objective function, we also desire policy $\pi$ to have the following properties:
  1. 1.
    Inductive: Policy $\pi$ is inductive if the parameters of the policy are independent of both the size of the graph and the number of partitions $k$. If the policy is not inductive then it will be unable to infer on unseen size graphs/ number of partitions.
  2. 2.
    Learning Versatile Objectives: To optimize the parameters of the policy, a target objective function is required. The optimization objective may not be differentiable and it might not be always possible to obtain a differentiable formulation. Hence, the policy $\pi$ should be capable of learning to optimize for a target objective that may or may not be differentiable.
Figure 1: Architecture of NeuroCUT. First the initial partitioning of the graph is performed based upon node’s initial embeddings. The initial embeddings are refined using GNN to infuse toplogical information from neighborhood. At each step a node is selected and its partitioned is updated. During the training phase the GNN parameters are updated and hence embeddings are re-computed. During inference, the GNN is called only once to compute the embeddings of the nodes of the graph.
Symbol Meaning
$\mathcal{G}$
Graph
$\mathcal{V}$
Node set
$e$
Edge $e\in\mathcal{E}$
$\mathcal{E}$
Edge set
$\mathcal{N}_{v}$
Neighboring nodes of node $v$
$Obj(\mathcal{G},\mathcal{P})$
Objective function based upon graph $\mathcal{G}$ and its partitioning $\mathcal{P}$
$k$
Number of partitions
$\mathcal{P}^{t}$
Partitioning at time $t$
${P}^{t}_{i}$
$i^{th}$ partition at time $t$
$\overline{{P}_{i}}$
Set of nodes that are not in the $i^{th}$ partition
$\mathcal{S}^{t}$
State of system at step $t$
$\operatorname{PART}(\mathcal{P}^{t},v)$
Partition of node $v$ at time $t$
$\mathcal{S}^{t}$
State representation of Partitions and Graph at step $t$
$\mathbf{pos}(v)$
Initial feature representation of node $v$
$\alpha$
Number of anchor nodes for lipschitz embedding
$T$
Length of trajectory
$\pi$
Policy function
Table 1: Notations used in the paper
3 NeuroCUT: Proposed Methodology
Fig. 1 describes the framework of NeuroCUT. For a given input graph $\mathcal{G}$, we first construct the initial partitions of nodes using a clustering based approach. Subsequently, a message-passing Gnn embeds the nodes of the graph ensuring inductivity to different graph sizes. Next, the assignment of nodes to partitions proceeds in a two-phased strategy. We first select a node to change its partition and then we choose a suitable partition for the selected node. Further, to ensure inductivity on the number of partitions, we decouple the number of partitions from the direct output representations of the model. This decoupling allows us to query the model to unseen number of partitions. After a node’s partition is updated, the reward with respect to change in partitioning objective value is computed and the parameters of the policy are optimized. The reward us learned through reinforcement learning (RL) (sutton2018reinforcement).
The choice of using RL is motivated two observations. Firstly, cut problems on graphs are generally recognized as NP-hard, making it impractical to rely on ground-truth data, which would be computationally infeasible to obtain. Secondly, the cut objective may lack differentiability. Therefore, it becomes essential to adopt a learning paradigm that can be trained even under these non-differentiable constraints. In this context, RL effectively addresses both of these critical requirements.
We now discuss each of the components in detail.
3.1 Learning Policy $\pi$ as Markov Decision Process
Given a graph $\mathcal{G}$, our objective is to find the partitioning $\mathcal{P}$ that maximizes/minimizes the target objective function $Obj(\mathcal{G},\mathcal{P})$. We model the task of iteratively updating the partition for a node as a Markov Decision Process (MDP) defined by the tuple $(S,\mathcal{A},\rho,R,\gamma)$. Here, $S$ is the state space, $\mathcal{A}$ is the set of all possible actions, $\rho:S\times S\times\mathcal{A}\rightarrow[0,1]$ denotes the state transition probability function, $R:S\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function and $\gamma\in(0,1)$ the discounting factor. We next formalize each of these notions in our MDP formulation.
Initialization. Instead of directly starting from empty partitions, we perform a warm start operation that clusters the nodes of the graph to obtain the initial graph partitions. The graph is clustered into $k$ clusters based upon their positional embeddings (discussed below). We use K-means (macqueen1967some) algorithm for this task. The details of the clustering are present in Appendix B.
Positional Embeddings: Given that the partitioning objectives are NP-hard mainly due to the combinatorial nature of the graph structure, we look for representations that capture the location of a node in the graph. Positional encodings provide an idea of the position in space of a given node within the graph. Two nodes that are closer in the graph, should be closer in the embedding space. Towards this, we use Lipschitz Embedding (bourgain1985lipschitz). Let $\mathcal{A}=\{a_{1},\cdots,a_{\alpha}\}\subseteq\mathcal{V}$ be a randomly selected subset of $\alpha$ nodes. We call them anchor nodes. From each anchor node $i$, the walker starts a random walk (bianchini2005inside) and jumps to a neighboring node $j$ with a transition probability ($\mathbf{W}_{ij}$) governed by the transition probability matrix $\mathbf{W}\in\mathbb{R}^{|V|\times|V|}$. Furthermore, at each step, with probability $c$ the walker jumps to a neighboring node $j$ and returns to the node $i$ with $1-c$. Let $\vec{r}_{ij}$ corresponds to the probability of the random walker starting from node $i$ and reaching node $j$.
$\vec{r}_{i}=c\tilde{\mathbf{W}}\vec{r}_{i}+(1-c)\vec{e}_{i}$ (6)
Eq. 6 describes the random walk starting at node $i$. In vector $\vec{e}_{i}\in\mathbb{R}^{|V|\times 1}$, only the $i^{th}$ element (the initial anchor node) is $1$, and the rest are set to zero. We set $\mathbf{W}_{ij}=\frac{1}{degree(j)}$ if edge $e_{ij}\in\mathcal{E}$, 00 otherwise. The random walk with restart process is repeated for $\beta$ iterations, where $\beta$ is a hyper-parameter. Here $\vec{e}_{i}\in\mathbb{R}^{|V|\times 1}$ and $c$ is a scalar.
Based upon the obtained random walk vectors for the set of anchor nodes $\mathcal{A}$, we embed all nodes $u\in\mathcal{V}$ in a $\alpha$-dimensional feature space:
$\mathbf{pos}(u)=[r_{1u},r_{2u},\cdots,r_{\alpha u}]$ (7)
State: The state space characterizes the state of the system at time $t$ in terms of the current set of partitions $\mathcal{P}^{t}$. Intuitively the state should contain information to help our model make a decision to select the next node and the partition for the node to be assigned. Let $\mathcal{P}^{t}$ denote the status of partitions at time $t$ wherein a partition $P^{t}_{i}$ is represented by all nodes belonging to $i^{th}$ partition. The state of the system at step $t$ is defined as
$S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{pos}(v)\;\;\forall v\in P_{i}^{t}\}\}$ (8)
Here state of each partition is represented by the collection of positional embedding of nodes in that partition.
Action: Towards finding the best partitioning scheme for the target partition objective we propose a 2-step action strategy to update the partitions of nodes. The first phase consists of identifying a node to update its partition. Instead of arbitrarily picking a node, we propose to prioritize selecting nodes for which the new assignment is more likely to improve the overall partitioning objective. In comparison to a strategy that arbitrarily selects nodes, the above mechanism promises greater improvement in the objective with less number of iterations. In the second phase, we calculate the score of each partition $\mathcal{P}$ with respect to the selected node from the first phase and then assign it to one of the partitions based upon the partitioning scores. We discuss both these phases in details below.
Phase 1: Node Selection Let $\operatorname{PART}(\mathcal{P}^{t},v)$ denote the partition of the node $v$ at step $t$. Our proposed formulation involves selecting a node $v$ at step $t$ belonging to partition $\operatorname{PART}(\mathcal{P}^{t},v)$ and then assigning it to a new partition. The newly assigned partition and the current partition of the node could also be same.
Towards this, we design a heuristic to prioritize selecting nodes which when placed in a new partition are more likely to improve the overall objective value. A node $v$ is highly likely to be moved from its current partition if most of its neighbours are in a different partition than that of node $v$. Towards this, we calculate the score of nodes $v\in\mathcal{V}$ in the graph as the ratio between the maximum number of neighbors in another partition and the number of neighbors in the same partition as $v$. Intuitively, if a partition exists in which the majority of neighboring nodes of a given node $v$ belong, and yet node $v$ is not included in that partition, then there is a high probability that node $v$ should be subjected to perturbation. Specifically the score of node $v$ at step $t$ is defined as:
$\operatorname{NodeScore}^{t}[v]=\\ {\frac{{\max_{p\in\mathcal{P}^{t}-\operatorname{PART}(\mathcal{P}^{t},v)}{|\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=p|}}}{|\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=\operatorname{PART}(\mathcal{P}^{t},v)|}}$ (9)
Further, a node having a higher degree implies it has several edges associated to it, hence, an incorrect placement of it could contribute to higher partitioning value. Hence, we normalize the scores by the $degree$ of the node.
Phase 2: Inductive Neural Method for Partition Selection Once a node is selected, the next phase involves choosing the new partition for the node. Towards this, we design an approach empowered by Graph Neural Networks (GNNs) which enables the model to be inductive with respect to size of graph. Further, instead of predicting a fixed-size score vector (nazi2019gap) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too. We discuss both above points in section below.
Message passing through graph neural network: To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (gcn; velivckovic2017graph; hamilton2017inductive). GNNs combine node feature information and the graph structure to learn better representations via feature propagation and aggregation.
We first initialize the input layer of each node $u\in\mathcal{V}$ in graph as $\mathbf{h}^{0}_{u}=\mathbf{pos}(u)$ using eq. 6. We perform $L$ layers of message passing to compute representations of nodes. To generate the embedding for node $u$ at layer $l+1$ we perform the following transformation(hamilton2017inductive):
$\mathbf{h}_{u}^{l+1}$ $=\mathbf{W}^{l}_{1}\mathbf{h}_{u}^{l}+\mathbf{W}^{l}_{2}\cdot\mathrm{mean}_{u^{\prime}\in\mathcal{N}_{u}}\mathbf{h}_{u^{\prime}}^{l}$ (10)
where $\mathbf{h}_{u}^{(l)}$ is the node embedding in layer $l$. $\mathbf{W}^{l}_{1}$ and $\mathbf{W}^{l}_{2}$ are trainable weight matrices at layer $l$.
Following $L$ layers of message passing, the final node representation of node $u$ in the $L^{th}$ layer is denoted by $\mathbf{h}_{u}^{L}\in\mathbb{R}^{d}$. Intuitively $\mathbf{h}_{u}^{L}$ characterizes $u$ using a combination of its own features and features aggregated from its neighborhood.
Scoring partitions: Recall from eq. 8, each partition at time $t$ is represented using the nodes belonging to that partition. Building upon this, we compute the score of each partition $p\in\mathcal{P}^{t}$ with respect to the node $v$ selected in Phase 1 using all the nodes in $p$. In contrast to predicting a fixed-size score vector corresponding to number of partitions (nazi2019gap), the proposed design choice makes the model inductive to the number of partitions. Specifically, the number of partitions are not directly tied to the output dimensions of the neural model.
Having obtained the transformed node embeddings through a Gnn in Eq. 10, we now compute the (unnormalized) score for node $v$ selected in phase 1 for each partition $p\in\mathcal{P}^{t}$ as follows:
$\operatorname{PartScore}$ $(p,v){=}\operatorname{AGG}(\{\operatorname{MLP}(\sigma(h_{v}|h_{u}))$ (11)
$\quad\quad\quad\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=p\})$
The above equation concatenates the selected node $v$’s embedding with its neighbors $u\in\mathcal{N}(v)$ that belong to the partition $p$ under consideration. In general, the strength of a partition assignment to a node is higher if its neighbors also belong to the same partition. The above formulation surfaces this strength in the embedding space. The concatenated representation $(h_{v}|h_{u})\forall u\in\mathcal{N}(v)$ is passed through an MLP that converts the vector into a score (scalar). We then apply an aggregation operator (e.g., mean) over all neighbors of $u$ belonging to $\mathcal{P}^{t}$ to get an unnormalized score for partition $p$. Here $\sigma$ is an activation function.
To compute the normalized score at step $t$ is finally calculated as softmax over the all partitions $p\in\mathcal{P}^{t}$ for the currently selected node $v$. Mathematically, the probability of taking action $a^{t}{=}p$ at time step $t$ at state $S^{t}$ is defined as:
$\pi((a^{t}{=}p)/S^{t})=\frac{exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$ (12)
During the course of trajectory of length $T$, we sample action $a^{t}\in\mathbb{R}$ i.e., the assignment of the partition for the node selected in phase 1 at step $t$ using policy $\pi$.
State Transition: After action $a^{t}$ is applied at state $S^{t}$, the state is updated to $S^{t+1}$ that involves updating the partition set $\mathcal{P}^{t+1}$. Specifically, if node $v$ belonged to $i^{th}$ partition at time $t$ and its partition has been changed to $j$ in phase 2, then we apply the below operations in order.
$P^{t+1}_{i}\leftarrow P^{t}_{i}\setminus v\text{ and }P^{t+1}_{j}\leftarrow P^{t}_{j}\cup v$ (13)
Reward: Our aim is to improve the value of the overall partitioning objective. One way is to define the reward $R^{t}$ at step $t\geq 0$ as the change in objective value of the partitioning i.e $Obj(\mathcal{G},\mathcal{P}^{t})$ at step $t$, i.e., $R^{t}=(Obj(\mathcal{G},\mathcal{P}^{t+1})-Obj(\mathcal{G},\mathcal{P}^{t}))/(Obj(\mathcal{G},\mathcal{P}^{t+1})+Obj(\mathcal{G},\mathcal{P}^{t}))$. Here, the denominator term ensures that the model receives a significant reward for even slight enhancements when dealing with a small objective value. However, this definition of reward focuses on short-term improvements instead of long-term. Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term. The discounted rewards are computed as the sum of the rewards over a horizon of actions with varying degrees of importance (short-term and long-term). Mathematically,
$D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}\vspace{-0.1in}$ (14)
where $T$ is the length of the horizon and $\gamma\in(0,1]$ is a discounting factor (hyper-parameter) describing how much we favor immediate rewards over the long-term future rewards.
The above reward mechanism provides flexibility to our framework to be versatile to objectives of different nature, that may or may not be differentiable. This is an advantage over existing neural method (nazi2019gap) where having a differentiable form of the partitioning objective is a pre-requisite.
3.2 Policy Loss Computation and Parameter Update
Our objective is to learn parameters of our policy network in such a way that actions that lead to an overall improvement of the partitioning objective are favored more over others. Towards this, we use REINFORCE gradient estimator (williams1992simple) to optimize the parameters of our policy network. Specifically, we wish to maximize the reward obtained for the horizon of length $T$ with discounted rewards $D^{t}$. Towards this end, we define a reward function $J(\pi_{\theta})$ as:
$J(\pi_{\theta})=\mathbb{E}\big{[}\sum_{t=0}^{T}\left(D^{t}\right)\big{]}$ (15)
We, then, optimize $J(\pi_{\theta})$ as follows:
${\nabla J(\pi_{\theta})=\left[\sum_{t=0}^{T}\left(D^{t}\right)\nabla_{\theta}log\pi_{\theta}({a}^{t}/S^{t})\right]}$ (16)
3.3 Training and Inference
For a given set of training graphs, we optimize the parameters of the policy network $\pi_{\theta}$ for $T$ steps. Subsequently, given a new graph and partition objective, we adapt the policy parameters $\pi_{\theta}$ using Eq. 16. Note that the trajectory length $T$ is not kept very large to avoid the long-horizon problem.
During inference, we compute the initial node embeddings, obtain initial partitioning and then run the forward pass of our policy to improve the partitioning objective over time.
3.4 Time complexity
The time complexity of NeuroCUT during inference is
$\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{E}|+k)\times T^{\prime}\big{)}$. Here $\alpha$ is number of anchor nodes, $\beta$ is number of random walk iterations, $k$ is the number of partitions and $T^{\prime}$ is the number of iterations during inference. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Further, for sparse graphs $\mathcal{E}=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$. In comparison to spectral clustering, which is a common graph partitioning algorithm having time complexity of $\mathcal{O}(|\mathcal{V}|^{3})$, the proposed technique’s complexity is much lower. This can also be attributed to the fact that NeuroCUT exploits the sparse nature of graphs unlike spectral clustering that performs matrix operations of $\mathcal{O}(|\mathcal{V}|^{3})$. For detailed derivation please see appendix  C.
[…]
Figure 2: Results of all methods on five datasets through radar plots. The color shade represents identity of each method. The legend in the first subplot is common for all plots. All values (for each partitioning objective) are normalized by dividing by the worst (maximum) cut value obtained on that dataset. The smaller the value, i.e, closer to the center, better is the performance. Further, when two polygons are overlapping, but not fully contained within each other, in general, the smaller polygon tends to indicate a better holistic performance across all objectives. Fig (a) represents cut values averaged across five datasets. Fig (b-f) depict results on each of the five datasets. The neural state-of-the-art GAP method produces the most inferior performance among considered baselines. Further in several cases it generates ”NaN” values (error). Our method NeuroCUT  either outperforms the baselines or produces the second best results while being close to the best one in all cases. Moreover, NeuroCUT  is the most robust and does not produce the worst cut values. The neural methods are trained and tested with the same number partitions ($k=5$). Unnormalized results for each dataset in table format are presented in Table  7 in appendix.
4 Experiments
In this section, we demonstrate the efficacy of NeuroCUT against state-of-the-art methods and establish that:
  • Efficacy and robustness: NeuroCUT consistently produces the best or the second best results in terms of the average performance over all objective functions. This establishes the robustness of NeuroCUT.
  • Inductivity: As one of the major strengths, unlike the state-of-the-art neural model GAP (nazi2019gap), NeuroCUT is inductive on the number of partitions and, consequently, generalizes well over unseen number of partitions.
4.1 Experimental Setup
Datasets:
We use five datasets for our experiments among which one is synthetic. They are as follows:
  • Cora and Citeseer (sen2008collective): These are citation networks. In these, a $k$-way cut could represent the partitioning of publications into $k$ clusters while minimizing the citations across the clusters. Each cluster might correspond to a distinct research field, shedding light on interdisciplinary connections or isolating specialized domains.
  • Facebook (leskovec2012learning): In social network analysis, a $k$-way cut can delineate communities of users with minimal inter-community connections. This partitioning unveils hidden social circles and provides insights into how different groups interact within the larger network. The datasets have multiple connected components, we work with the largest connected component to make sure we can evaluate proper cuts.
  • Harbin (li2019learning): In this road network, a $k$-way cut could entail partitioning road segments into $k$ subsets to optimize traffic flow between them. This partitioning aids in urban planning by identifying key intersections and routes that minimize congestion.
  • Stochastic Block Models (SBM) (abbe2017community): This synthetic network is generated by SBM where $k$-way cuts reveal the underlying community structure. This helps to validate our method by comparing the identified partitions with the ground truth communities.
Baselines:
We compare our proposed method NeuroCUT with the state-of-the-art neural method called GAP (nazi2019gap). It allows balanced normalized cuts with an end-to-end framework with a differentiable loss function which is a continuous relaxation version of normalized cut. We also use two popular non-neural baselines: hMetis (karypis1997multilevel) and spectral clustering (ng2001spectral).
Other settings:
We run all our experiments on an Ubuntu 20.04 system running on Intel Xeon 6248 processor with 96 cores and 1 NVIDIA A100 GPU with 40GB memory for our experiments. We used GraphSage(hamilton2017inductive) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$. We used Adam optimizer for training the parameters of our policy network $\pi_{\theta}$. For computing discounted reward in RL, we use discount factor $\gamma=0.99$. We set the length of trajectory $T$ during training as $2$. At time step $t$, the rewards are computed from time $t$ to $t+T$ and parameters of the policy $\pi_{\theta}$ are updated. The default number of anchor nodes for computing positional embeddings is set to $35$.
4.2 Results on Transductive Setting
We compare our proposed method NeuroCUT against all the baselines where the neural models are trained and tested with the same number of partitions ($k=5$). Fig. 2 presents the results for all datasets and all methods. The results are depicted using spider plots where the smaller the value i.e closer to the center, better is the performance. Further, in the plots, if two polygons are overlapping, but not fully contained within each other, in general the smaller polygon tends to indicate a better holistic performance across all objectives (heckert2002handbook). NeuroCUT outperforms the baselines or produces the next to optimal results. Moreover, NeuroCUT is the most robust among all the methods and does not produce the worst cut values in any of the settings. Overall, the performance of NeuroCUT is the most consistent among all the methods. Unnormalized results for each dataset in table format are presented in Table  7 in appendix.
NeuroCUT incorporates causality during inference and this leads to the robust performance observed in fig. 2. Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in GAP (nazi2019gap). We observe that the non-neural methods fail to generalize well across all objective functions since they do not learn the base heuristic as a function of the objective and the dataset. Finally, we note that in SBMs the task is easier as there are already $5$ clusters which aligns well with the the desired number of partitions, which is also $5$.
4.3 Results on Inductive Setting
Dataset Method $normalised$ $sparsest$ $balanced$ $kmin$
Cora hMetis 0.772 0.772 3.177 0.097
Spectral 0.723 1.503 1.845 0.008
NeuroCUT 0.591 2.451 0.981 0.022
CiteSeer hMetis 0.415 0.415 1.554 0.056
Spectral 0.328 1.094 0.792 0.003
NeuroCUT 0.363 1.060 0.457 0.031
SBMs hMetis 3.259 3.260 65.955 0.393
Spectral 2.495 2.513 50.805 0.242
NeuroCUT 3.241 5.505 45.602 0.038
Table 2: Inductive results based on the number of partitions of all methods on all six datasets. The best ones are shown by green and the second best ones are in gray. Lower values are better. Our method NeuroCUT either outperforms the baselines or produces the second best results while being close to the best one in ten out of twelve settings. NeuroCUT is trained on different partition sizes ($k=5$ and $k=10$) and the test data involves an unseen partition size ($k=8$). GAP is not inductive in terms of number of partitions and thus the results are not present here.
We show that NeuroCUT generalizes to unknown number of partitions. We compare it against only the non-neural baselines since GAP does not generalize to unseen number of partitions. Here, NeuroCUT is trained on different partition sizes ($k=5$ and $k=10$) and the test data involves an unseen partition size ($k=8$). Table 2 shows the results. Our method NeuroCUT either outperforms the baselines or produces the second best results while being close to the best one in most of the settings. Note that while the non-neural methods need to be re-run for the unseen partitions, NeuroCUT only runs forward pass to produce the results on an unseen partition size.
4.4 Generalization to Topological changes in graphs
In this section, we analyze the performance of NeuroCUT to topological changes in the graph. Specifically, once a model is trained, we aim to evaluate its performance when the topology of the graph changes which is a common scenario in dynamic networks such as social networks where connections are formed/deleted over time.
Towards this first we train NeuroCUT on a given graph. Next, we add/remove $x\%$ percentage of edges in the graph. Then we predict the cuts on the perturbed graph. To predict cuts for the perturbed graph, our model NeuroCUT only needs to just run a forward pass over it.
Tables  3,4,5,6 show the results on two datasets, namely, Cora and Facebook for different partitioning objectives at different degrees of perturbations. In almost all settings, our method NeuroCUT outperforms the baselines. The only neural baseline i.e GAP had to be retrained on perturbed graph and still produced NaN values(indicated by -) in most of the cases. These results show that our proposed method NeuroCUT is more robust in comparison to existing neural methods at different levels of perturbations. Additionally, in several cases it outperforms non-neural methods which were directly optimized on perturbed graphs.
[…]
Cora hMetis No 0.351 0.355 0.399 0.435 0.506
Spectral No 0.346 0.442 0.479 0.424 0.490
GAP Yes 0.512 0.500 0.483 0.808 0.664
NeuroCUT Yes 0.321 0.307 0.364 0.428 0.569
Facebook hMetis No 1.112 1.113 1.126 1.126 1.131
Spectral No 2.034 1.676 1.918 1.92 1.679
GAP Yes - - 0.482 - -
NeuroCUT Yes 0.487 0.459 0.463 0.494 0.864
Table 3: Perturbing $x\%$ Edges - Normalised Cuts
[…]
Cora hMetis No 1.442 1.464 1.649 1.819 2.169
Spectral No 0.795 1.058 1.224 0.957 1.452
GAP Yes 2.029 1.978 1.915 3.167 2.816
NeuroCUT Yes 1.015 1.317 1.479 1.546 2.053
Facebook hMetis No 51.669 52.937 52.110 52.962 52.236
Spectral No 5.625 4.722 5.591 5.591 4.867
GAP Yes - - 11.886 - -
NeuroCUT Yes 3.528 3.516 3.745 4.060 5.298
Table 4: Perturbing $x\%$ Edges - Sparsest Cuts
[…]
Cora hMetis No 0.352 0.355 0.399 0.435 0.506
Spectral No 1.089 1.223 0.800 1.207 0.841
GAP Yes 0.536 0.529 0.515 0.870 0.681
NeuroCUT Yes 0.526 0.544 0.613 0.651 0.836
Facebook hMetis No 1.112 1.113 1.126 1.127 1.131
Spectral No 2.818 2.459 2.697 2.697 2.456
GAP Yes - - 0.506 - -
NeuroCUT Yes 0.831 0.793 0.800 0.855 0.921
Table 5: Perturbing $x\%$ Edges - Balanced Cuts
[…]
Cora hMetis No 0.071 0.072 0.080 0.087 0.101
Spectral No 0.002 0.001 0.054 0.001 0.064
GAP Yes 0.099 0.097 0.094 0.114 0.126
NeuroCUT Yes 0.010 0.013 0.014 0.018 0.021
Facebook hMetis No 0.199 0.203 0.201 0.203 0.201
Spectral No 0.000 0.000 0.000 0.000 0.000
GAP Yes 0.024 0.026 0.044 0.026 0.027
NeuroCUT Yes 0.000 0.000 0.000 0.000 0.000
Table 6: Perturbing $x\%$ Edges - KMin Cuts
[…]
Figure 3: Results on the initial and the final cut values at $k=5$. It shows that our neural model NeuroCUT (Final) performs more accurate node and partition selection to optimize the objective function. Subsequently, there is a significant difference between the initial and final cut values.
4.5 Ablation Studies
Initial vs Final Cut Values. NeuroCUT takes a warm start by partitioning nodes based on clustering over their Lipschitz embeddings, which are subsequently fine-tuned auto-regressively through the reinforcement learning based framework. How much does the partitioning objective improve since the initial clustering? Figure 3 sheds light on this question. Specifically, it shows the difference between initial cut value due after clustering and the final cut value from the partitions produced by our method. We note that there is a significant difference between the initial and final cut values, which essentially shows the effectiveness of NeuroCUT.
[…]
Figure 4: NeuroCUT results with heuristics for node selection in Phase 1 for $k=5$. Our heuristic finds better cuts than a random selection algorithm for two objective functions: Normalized Cut and KMin Cut.
Impact of Node Selection Procedures. Figure 4 shows how our proposed heuristic for node selection helps in finding the near-optimal cuts for any given objective function. We compare and show that a simpler node selection where we select all the nodes one by one in a random order produces substantially inferior results. Furthermore, the difference lies not only in final cut-values but also we require much less perturbations.
In App. D of appendix, we present additional results on variation of parameters of NeuroCUT.
5 Conclusion
In this paper, we have introduced a novel graph reinforcement learning framework NeuroCUT to effectively address the graph partitioning problem. A distinctive feature of our framework is its adaptability to diverse partitioning objectives that may or may not be differentiable, which sets it apart from conventional algorithms. Our approach has also shown the ability to generalize on unseen graphs as well as unseen number of partitions. The efficacy of our approach is empirically validated through an extensive evaluation on six datasets with four graph partitioning objectives. Notably, our method shows remarkable performance when compared to both non-neural and neural state-of-the-art techniques, proving its competence in both inductive and transductive settings. In summary, NeuroCUT offers a versatile, inductive, and empirically validated solution for graph partitioning.
Dataset $|\mathcal{V}|$ $|\mathcal{E}|$ Method $normalised$ $sparsest$ $balanced$ $kmin$ AVG DIFF MAX DIFF
ACTDIFF ACTDIFF ACTDIFF ACTDIFF
Cora 2485 5069 hMetis 0.343   0.115 1.412   0.061 0.345   0.000 0.069   0.063 0.060 0.115
Spectral 0.370   0.142 1.351   0.000 1.165   0.820 0.006   0.000 0.240 0.820
GAP 0.505   0.277 2.074   0.723 0.536   0.191 0.095   0.089 0.320 0.723
NeuroCUT 0.228   0.000 1.351   0.000 0.570   0.225 0.011   0.005 0.057 0.225
CiteSeer 2120 3679 hMetis 0.198   0.025 0.673   0.227 0.198   0.000 0.038   0.037 0.072 0.227
Spectral 0.173   0.000 0.446   0.000 0.929   0.731 0.001   0.000 0.182 0.731
GAP 0.267   0.094 0.795   0.349 0.312   0.114 0.043   0.042 0.150 0.349
NeuroCUT 0.211   0.038 0.712   0.266 0.296   0.098 0.036   0.035 0.109 0.266
SBMs 500 5150 hMetis 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
Spectral 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
GAP - - - 0.035   $0$
NeuroCUT 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
Facebook 1034 26749 hMetis 1.115   0.858 52.319   48.198 1.115   0.143 0.200   0.197 12.349 48.198
Spectral 1.670   1.413 4.722   0.601 2.459   1.487 0.003   0.000 0.875 1.487
GAP - - - 0.025   0.022 - -
NeuroCUT 0.257   0.000 4.121   0.000 0.972   0.000 0.015   0.012 0.003 0.012
Harbin 6235 15205 hMetis 0.058 0.000 0.184   0.000 0.058   0.000 0.011   0.000 0.000 0.000
Spectral 0.104   0.046 0.348   0.164 0.403   0.345 0.015   0.004 0.140 0.345
GAP 0.175   0.117 0.551   0.367 0.269   0.211 0.017   0.006 0.175 0.367
NeuroCUT 0.092   0.034 0.286   0.102 0.099   0.041 0.018   0.007 0.046 0.102
Table 7: Results of all methods on all five datasets. “ACT” denotes the actual values of the measures and “DIFF” shows the difference with the best value for the dataset. “AVG DIFF” and “MAX DIFF” present the average and the maximum of the differences over all the measures respectively. The best average differences are shown by green and the second best ones are in gray. The worst result, i.e., the maximum difference is shown in red. The neural state-of-the-art GAP method produces the most inferior performance among considered baselines. Our method NeuroCUT either outperforms the baselines (green) or produces the second best (gray) results while being close to the best one in all cases. Moreover, NeuroCUT is the most robust and never produces the worst maximum difference (i.e., in red). The neural methods are trained and tested with the same number partitions ($k=5$).
Appendix A Result for all datasets in transductive setting
In Table  7 we present the results obtained by all methods on different datasets. These correspond to radar plot obtained in fig. 2 in main paper .
Appendix B Clustering for Initialization
As discussed in sec 3.1 in main paper, we first cluster the nodes of the graph into $k$ clusters where $k$ is the number of partitions. Towards this, we apply K-means algorithm (macqueen1967some) on the nodes of the graph where a node is represented by its positional node embedding i.e $pos(u)\;\forall u\in\mathcal{V}$. Further, we used $L_{\inf}$ norm as the distance metric for clustering.
Appendix C Time complexity of NeuroCUT
We derive the time complexity of forward pass of NeuroCUTas discussed in sec. 3.4 in main paper.
(1) First the positional embeddings for all nodes in the graph are computed. This involves running RWR for $\alpha$ anchor nodes for $\beta$ iterations(Eq.  6). This takes $\mathcal{O}(\alpha\times\beta)$ time. (2)Next, GNN is called to compute embeddings of node. In each layer of GNN a node $v\in\mathcal{V}$ aggregates message from $d$ neighbors where $d$ is the average degree of a node. This takes $\mathcal{O}(\mathcal{V})$ time. This operation is repeat for $L$ layers. Since $L$ is typically 1 or 2 hence we ignore this factor. (3) The node selection algorithm is used that computes score for each node based upon its neighborhood using eq. 9. This consumes $\mathcal{O}(|\mathcal{V}|\times d)$ time. (4) Finally for the selected node, its partition has to be determined using eq. 11 and  12. This takes $\mathcal{O}(d+k)$ time, as we consider only the neighbors of the selected node to compute partition score. Steps 2-4 are repeated for $T^{\prime}$ iterations. Hence overall running time is $\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{V}|\times d+k)\times T^{\prime}\big{)}$. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Since $|\mathcal{V}|\times d\approx|\mathcal{E}|$, hence complexity is $o(|\mathcal{E}|\times|\mathcal{V}|\big{)}$ Further, for sparse graphs $|\mathcal{E}|=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.
Appendix D Impact of number of anchor nodes
In this section we analyze the impact of number of anchor nodes chosen for computing positional embeddings of nodes in the graph $pos(u)$. Fig. 5 shows the results where $Y$-axis shows the normalized cut value and $X$-axis shows the number of anchor nodes. The best results are achieved at the number of anchor nodes being 35 except in Citeseer. The other variation in the number of anchor nodes produces similar results often.
Figure 5: Results on varying the number of anchor nodes on the normalized cut value produced by NeuroCUT.

### v2

@@
 The majority of formulations related to graph partitioning exhibit NP-hardness due to their combinatorial nature.
-As a result, conventional approximation algorithms rely on heuristic methods, sometimes with approximation guarantees and sometimes without.
-Unfortunately, traditional approaches are tailored for specific partitioning objectives and do not generalize well across other known partitioning objectives from the literature.
-To overcome this limitation, and learn heuristics from the data directly, neural approaches have emerged, demonstrating promising outcomes.
-In this study, we extend this line of work through a novel framework, NeuroCUT.
-NeuroCUT introduces two key innovations over prevailing methodologies.
-First, it is inductive to both graph topology and the partition count, which is provided at query time.
-Second, by leveraging a reinforcement learning based framework over node representations derived from a graph neural network, NeuroCUT can accommodate any optimization objective, even those encompassing non-differentiable functions.
-Through empirical evaluation, we demonstrate that NeuroCUT excels in identifying high-quality partitions, showcases strong generalization across a wide spectrum of partitioning objectives, and exhibits resilience to topological modifications.
+Conventional methods, like approximation algorithms or heuristics, are designed for distinct partitioning objectives and fail to achieve generalization across other important partitioning objectives.
+Recently machine learning-based methods have been developed that learn directly from data.
+Further, these methods have a distinct advantage of utilizing node features that carry additional information.
+However, these methods assume differentiability of target partitioning objective functions and cannot generalize for an unknown number of partitions, i.e., they assume the number of partitions is provided in advance.
+In this study, we develop NeuroCUT with two key innovations over previous methodologies.
+First, by leveraging a reinforcement learning-based framework over node representations derived from a graph neural network and positional features, NeuroCUT can accommodate any optimization objective, even those with non-differentiable functions.
+Second, we decouple the parameter space and the partition count making NeuroCUT inductive to any unseen number of partition, which is provided at query time.
+Through empirical evaluation, we demonstrate that NeuroCUT excels in identifying high-quality partitions, showcases strong generalization across a wide spectrum of partitioning objectives, and exhibits strong generalization to unseen partition count.
 Abstract
@@
 The majority of the formulations are NP-hard and thus the proposed solutions are either heuristics or algorithms with approximate solutions (karypis1997multilevel).
-In recent years there has been a growing interest in using machine learning techniques to design heuristics for NP-Hard combinatorial problems on graphs.
-This line of work is motivated from the observation that traditional algorithmic approaches are limited to using the same heuristic irrespective of the underlying data distribution.
-Neural approaches, instead, focus on learning the heuristic as a function of the data without deriving explicit algorithms (duong2023deep; jung2022learning; nazi2019gap).
-Significant advances have been made in learning to solve a variety of combinatorial optimization problems such as Set Cover, Travelling Salesman Problem, Influence Maximization etc.
-(khalil2017learning; gcomb; kool2018attention; joshi2019efficient).
-The neural architectural designs in these methods are generally suitable for a specific set of CO problems (mazyavkina2021reinforcement) of similar nature.
-For instance, AM  (kool2018attention) and NeuroLKH (xin2021neurolkh) are targeted towards routing problems.
-GCOMB (gcomb) and LeNSE (ireland2022lense) aim to solve problems that are submodular in nature.
-Owing to these assumptions, the architectural designs in these works are not suitable for our problem of $k$-way inductive graph partitioning with versatile partitioning objectives.
-The most relevant method to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (nazi2019gap).
+Recently, there have been attempts to solve graph partitioning problem via neural approaches.
+The neural approaches have a distinct advantage that they can utilize node features.
+Node features supply additional information and provide contextual insights that may improve the accuracy of graph partitioning.
+For instance, (wang2017mgae) has recognized the importance of incorporating node attributes in graph clustering where nodes are partitioned into disjoint groups.
+Note that there are existing neural approaches (khalil2017learning; gcomb; kool2018attention; joshi2019efficient) to solve other NP-hard graph combinatorial problems (e.g., minimum vertex cover).
+However these methods are not generic enough to solve graph partitioning.
+In this paper, we build a single framework to solve several graph partitioning problems.
+One of the most relevant to our work is the method DMoN by muller2023graph.
+This method designs a neural architecture for cluster assignments and use a modularity-based objective function for optimizing these assignments.
+Another method, that is relevant to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (nazi2019gap).
 It proposes a differentiable loss function for partitioning based on a continuous relaxation of the normalized cut formulation.
@@
 The idea is to construct a reformulation of the multicut ILP constraints to a polynomial program as a loss function.
+However, the problem formulation is different than the generic normalized cut or mincut problem.
 Finally,  (gatti2022graph) solves the normalized cut problem only for the case where the number of partitions is exactly two.
-Nevertheless, existing neural approaches that tackle graph partitioning problem suffer three notable drawbacks, which we address.
-  • Assumption of differential objective function: Existing neural approaches assume the objective function to be differentiable.
-As we will illustrate in § 2, the assumption does not always hold in the real-world.
-  • Inductivity to partition count: The number of partitions required to segment a graph is an input parameter.
+Nevertheless, these neural approaches for the graph partitioning problem often do not use node features and only limited to a distinct partitioning objective.
+Here, we point out notable drawbacks that we address in our framework.
+  • Non-inductivity to partition count: The number of partitions required to segment a graph is an input parameter.
 Hence, it is important for a learned model to generalize to any partition count without retraining.
 Existing neural approaches are non-inductive to the number of partitions, i.e they can only infer on number of partitions on which they are trained.
-Additionally, it’s worth noting that the optimal number of partitions is often unknown beforehand.
+Additionally, it is worth noting that the optimal number of partitions is often unknown beforehand.
 Hence, it is a common practice to experiment with different partition counts and evaluate their impact on the partitioning objective.
-Typically, the optimal partition count corresponds to the point where the second-order derivative of the objective function approaches saturation.
-Therefore, inductivity to the number of partitions becomes an essential requirement for conducting such experiments in a scalable manner.
-  • Generalizability to objective functions: Multiple objective functions have been studied in the partitioning literature.
+While the existing methods (muller2023graph; nazi2019gap; gatti2022graph) assume that the number of partitions ($k$) is known beforehand, our proposed method can generalize to any $k$.
+  • Non-generalizability to different cut functions: Multiple objective functions for graph cut have been studied in the partitioning literature.
 The optimal objective function hinges upon the subsequent application in question.
-Consequently, it holds significance for an algorithmic framework to possess the capacity to encompass a diverse array of objective functions.
-Unfortunately, prevailing neural frameworks are founded upon presumptions that yield efficacy exclusively within the confines of their designated objective function.
+For instance, the two most relevant studies, DMoN (muller2023graph) and GAP (nazi2019gap) focus on maximizing modularity and minimizing normalized cut respectively.
+Our framework is generic and can solve different partitioning objectives.
+  • Assumption of differential objective function: Existing neural approaches assume the objective function to be differentiable.
+As we illustrate in § 2, the assumption does not always hold in the real-world.
+For instance, the sparsest cut (chawla2006hardness) and balanced cut (nazi2019gap) objectives are not differentiable.
 1.1 Contributions:
 In this paper, we circumvent the above-mentioned limitations through the following key contributions.
-  • Versatile objectives: We develop an auto-regressive, graph reinforcement learning framework, called NeuroCUT, to solve the graph partitioning problem.
-Diverging from conventional algorithms, NeuroCUT can handle different partitioning objectives.
+  • Versatile objectives: We develop NeuroCUT; an auto-regressive, graph reinforcement learning framework that integrates positional information, to solve the graph partitioning problem for attributed graphs.
+Diverging from conventional algorithms, NeuroCUT can solve multiple partitioning objectives.
 Moreover, unlike other neural methods, NeuroCUT can accommodate diverse partitioning objectives, without the necessity for differentiability.
-  • Inductivity to topology and number of partitions: The parameter space of NeuroCUT is independent of the graph size or the partition count.
-This innovative architecture endows NeuroCUT with the ability to generalize effectively to previously unseen graphs of varying scales and adapt to arbitrary partition count specified during inference.
+  • Inductivity to number of partitions: The parameter space of NeuroCUT is independent of the partition count.
+This innovative decoupled architecture endows NeuroCUT with the ability to generalize effectively to arbitrary partition count specified during inference.
   • Empirical Assessment: We perform comprehensive experiments employing real-world datasets, evaluating NeuroCUT across four distinct graph partitioning objectives.
-Our empirical investigation substantiates the efficacy of NeuroCUT in partitioning tasks, showcasing its robustness across a spectrum of objective functions, and demonstrating its capacity to generalize effectively to partition sizes and topological configurations that it has not encountered during training.
+Our empirical investigation substantiates the efficacy of NeuroCUT in partitioning tasks, showcasing its robustness across a spectrum of objective functions.
+We also demonstrate the capacity of NeuroCUT to generalize effectively to partition sizes that it has not encountered during training.
 2 Problem Formulation
-In this section, we introduce the concepts central to our work and formulate the problem. All the notations used in this work are outlined in Tab. 1 in appendix.
+In this section, we introduce the concepts central to our work and formulate the problem. All the notations used in this work are outlined in Table 1.
 Definition 1 (Graph)
-We denote a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{G}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges.
+We denote a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E},\textbf{{X}})$ where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges and $\textbf{{X}}\in\mathbb{R}^{|\mathcal{V}|\times|F|}$ refers to node feature matrix where $F$ is the set of all features in graph $\mathcal{G}$.
 Definition 2 (Cut)
@@
   2. 2.
-    Normalized Mincut (shi2000normalized): The $k$-minimum cut criteria favors cutting small sets of isolated nodes in the graph.
+    Normalized Cut (shi2000normalized): The $k$-MinCut criteria favors cutting small sets of isolated nodes in the graph.
 To avoid this unnatural bias for partitioning out small sets of points, normalised cut computes the cut cost as a fraction of the total edge connections to all the nodes in the graph.
@@
   4. 4.
-    (Weighted) Sparsest Cuts (chawla2006hardness): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition.
+    Sparsest Cuts (chawla2006hardness): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition.
 We generalize it to $k$-way sparsest cuts by summing up the value for all the partitions.
@@
 Figure 1: Architecture of NeuroCUT.
-First the initial partitioning of the graph is performed based upon node’s initial embeddings.
-The initial embeddings are refined using GNN to infuse toplogical information from neighborhood.
+First, the initial partitioning of the graph is performed based on node raw features and positional embeddings.
+These embeddings are refined using GNN to infuse toplogical information from neighborhood.
 At each step a node is selected and its partitioned is updated.
@@
 Edge set
+X
+Feature matrix containing raw node features
 $\mathcal{N}_{v}$
@@
 $\mathbf{pos}(v)$
-Initial feature representation of node $v$
+Positional embedding of node $v$
+$\mathbf{emb_{init}}(v)$
+Initial embedding of node $v$
 $\alpha$
@@
 Further, to ensure inductivity on the number of partitions, we decouple the number of partitions from the direct output representations of the model.
-This decoupling allows us to query the model to unseen number of partitions.
+This decoupling allows us to query the model to an unseen number of partitions.
 After a node’s partition is updated, the reward with respect to change in partitioning objective value is computed and the parameters of the policy are optimized.
-The reward us learned through reinforcement learning (RL) (sutton2018reinforcement).
-The choice of using RL is motivated two observations.
+The reward is learned through reinforcement learning (RL) (sutton2018reinforcement).
+The choice of using RL is motivated through two observations.
 Firstly, cut problems on graphs are generally recognized as NP-hard, making it impractical to rely on ground-truth data, which would be computationally infeasible to obtain.
@@
 In this context, RL effectively addresses both of these critical requirements.
-We now discuss each of the components in detail.
-3.1 Learning Policy $\pi$ as Markov Decision Process
+Additionally, in the process of sequentially constructing a solution, RL allows us to model the gain obtained by perturbing the partition of a node.
+Markov Decision Process.
 Given a graph $\mathcal{G}$, our objective is to find the partitioning $\mathcal{P}$ that maximizes/minimizes the target objective function $Obj(\mathcal{G},\mathcal{P})$.
@@
 We next formalize each of these notions in our MDP formulation.
+3.1 State: Initialization & Positional Encoding
 Initialization.
 Instead of directly starting from empty partitions, we perform a warm start operation that clusters the nodes of the graph to obtain the initial graph partitions.
-The graph is clustered into $k$ clusters based upon their positional embeddings (discussed below).
+The graph is clustered into $k$ clusters based on their raw features and positional embeddings (discussed below).
 We use K-means (macqueen1967some) algorithm for this task.
-The details of the clustering are present in Appendix B.
-Positional Embeddings: Given that the partitioning objectives are NP-hard mainly due to the combinatorial nature of the graph structure, we look for representations that capture the location of a node in the graph.
+The details of the clustering are present in Appendix A.2.
+Positional Encoding (Embeddings).
+Given that the partitioning objectives are NP-hard mainly due to the combinatorial nature of the graph structure, we look for representations that capture the location of a node in the graph.
 Positional encodings provide an idea of the position in space of a given node within the graph.
@@
 $\mathbf{pos}(u)=[r_{1u},r_{2u},\cdots,r_{\alpha u}]$ (7)
-State: The state space characterizes the state of the system at time $t$ in terms of the current set of partitions $\mathcal{P}^{t}$.
+To accommodate both raw feature and positional information we concatenate the raw node features i.e $\textbf{X}[u]$ with the positional embedding $\mathbf{pos}(u)$ for each node $u$ to obtain the initial embedding which will act as input to our neural model.
+Specifically,
+$\mathbf{emb_{init}}(u)=\textbf{X}[u]\;\;\mathbin{\|}\;\;\mathbf{pos}(u)$ (8)
+[…]
+State.
+The state space characterizes the state of the system at time $t$ in terms of the current set of partitions $\mathcal{P}^{t}$.
 Intuitively the state should contain information to help our model make a decision to select the next node and the partition for the node to be assigned.
@@
 The state of the system at step $t$ is defined as
-$S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{pos}(v)\;\;\forall v\in P_{i}^{t}\}\}$ (8)
-Here state of each partition is represented by the collection of positional embedding of nodes in that partition.
-Action: Towards finding the best partitioning scheme for the target partition objective we propose a 2-step action strategy to update the partitions of nodes.
+$S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{emb_{init}}(v)\;\;\forall v\in P_{i}^{t}\}\}$ (9)
+Here state of each partition is represented by the collection of initial embedding of nodes in that partition.
+3.2 Action: Selection of Nodes and Partitions
+Towards finding the best partitioning scheme for the target partition objective we propose a 2-step action strategy to update the partitions of nodes.
 The first phase consists of identifying a node to update its partition.
@@
 We discuss both these phases in details below.
-Phase 1: Node Selection Let $\operatorname{PART}(\mathcal{P}^{t},v)$ denote the partition of the node $v$ at step $t$.
+Phase 1: Node selection to identify node to perturb
+Let $\operatorname{PART}(\mathcal{P}^{t},v)$ denote the partition of the node $v$ at step $t$.
 Our proposed formulation involves selecting a node $v$ at step $t$ belonging to partition $\operatorname{PART}(\mathcal{P}^{t},v)$ and then assigning it to a new partition.
@@
 Specifically the score of node $v$ at step $t$ is defined as:
-$\operatorname{NodeScore}^{t}[v]=\\ {\frac{{\max_{p\in\mathcal{P}^{t}-\operatorname{PART}(\mathcal{P}^{t},v)}{|\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=p|}}}{|\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=\operatorname{PART}(\mathcal{P}^{t},v)|}}$ (9)
+$\operatorname{NodeScore}^{t}[v]=\\ {\frac{{\max_{p\in P^{t}-\operatorname{PART}(P^{t},v)}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=p|}}}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=\operatorname{PART}(P^{t},v)|}}$ (10)
 Further, a node having a higher degree implies it has several edges associated to it, hence, an incorrect placement of it could contribute to higher partitioning value.
 Hence, we normalize the scores by the $degree$ of the node.
-Phase 2: Inductive Neural Method for Partition Selection Once a node is selected, the next phase involves choosing the new partition for the node.
-Towards this, we design an approach empowered by Graph Neural Networks (GNNs) which enables the model to be inductive with respect to size of graph.
-Further, instead of predicting a fixed-size score vector (nazi2019gap) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too.
+Phase 2: Inductive Method for Partition Selection
+Once a node is selected, the next phase involves choosing the new partition for the node.
+Towards this, we design an approach empowered by Graph Neural Networks (GNNs) (hamilton2017inductive) which enables the model to be inductive with respect to size of graph.
+Further, instead of predicting a fixed-size score vector (nazi2019gap; tsitsulin2023graph) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too.
 We discuss both above points in section below.
-Message passing through graph neural network: To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (gcn; velivckovic2017graph; hamilton2017inductive).
+Message passing through Graph Neural Network
+To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (hamilton2017inductive).
 GNNs combine node feature information and the graph structure to learn better representations via feature propagation and aggregation.
-We first initialize the input layer of each node $u\in\mathcal{V}$ in graph as $\mathbf{h}^{0}_{u}=\mathbf{pos}(u)$ using eq.
-6.
+We first initialize the input layer of each node $u\in\mathcal{V}$ in graph as $\mathbf{h}^{0}_{u}=\mathbf{emb_{init}}(u)$ using eq.
+8.
 We perform $L$ layers of message passing to compute representations of nodes.
 To generate the embedding for node $u$ at layer $l+1$ we perform the following transformation(hamilton2017inductive):
-$\mathbf{h}_{u}^{l+1}$ $=\mathbf{W}^{l}_{1}\mathbf{h}_{u}^{l}+\mathbf{W}^{l}_{2}\cdot\mathrm{mean}_{u^{\prime}\in\mathcal{N}_{u}}\mathbf{h}_{u^{\prime}}^{l}$ (10)
+$\mathbf{h}_{u}^{l+1}$ $=\mathbf{W}^{l}_{1}\mathbf{h}_{u}^{l}+\mathbf{W}^{l}_{2}\cdot\frac{1}{|\mathcal{N}_{u}|}\sum_{u^{\prime}\in\mathcal{N}_{u}}\mathbf{h}_{u^{\prime}}^{l}$ (11)
 where $\mathbf{h}_{u}^{(l)}$ is the node embedding in layer $l$. $\mathbf{W}^{l}_{1}$ and $\mathbf{W}^{l}_{2}$ are trainable weight matrices at layer $l$.
@@
 Intuitively $\mathbf{h}_{u}^{L}$ characterizes $u$ using a combination of its own features and features aggregated from its neighborhood.
-Scoring partitions: Recall from eq.
-8, each partition at time $t$ is represented using the nodes belonging to that partition.
+Scoring partitions.
+Recall from eq.
+9, each partition at time $t$ is represented using the nodes belonging to that partition.
 Building upon this, we compute the score of each partition $p\in\mathcal{P}^{t}$ with respect to the node $v$ selected in Phase 1 using all the nodes in $p$.
-In contrast to predicting a fixed-size score vector corresponding to number of partitions (nazi2019gap), the proposed design choice makes the model inductive to the number of partitions.
+In contrast to predicting a fixed-size score vector corresponding to number of partitions (nazi2019gap; tsitsulin2023graph), the proposed design choice makes the model inductive to the number of partitions.
 Specifically, the number of partitions are not directly tied to the output dimensions of the neural model.
@@
 Mathematically, the probability of taking action $a^{t}{=}p$ at time step $t$ at state $S^{t}$ is defined as:
-$\pi((a^{t}{=}p)/S^{t})=\frac{exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$ (12)
+$\pi((a^{t}{=}p)/S^{t})=\frac{\exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$ (13)
 During the course of trajectory of length $T$, we sample action $a^{t}\in\mathbb{R}$ i.e., the assignment of the partition for the node selected in phase 1 at step $t$ using policy $\pi$.
-State Transition: After action $a^{t}$ is applied at state $S^{t}$, the state is updated to $S^{t+1}$ that involves updating the partition set $\mathcal{P}^{t+1}$.
+State Transition.
+After action $a^{t}$ is applied at state $S^{t}$, the state is updated to $S^{t+1}$ that involves updating the partition set $\mathcal{P}^{t+1}$.
 Specifically, if node $v$ belonged to $i^{th}$ partition at time $t$ and its partition has been changed to $j$ in phase 2, then we apply the below operations in order.
 $P^{t+1}_{i}\leftarrow P^{t}_{i}\setminus v\text{ and }P^{t+1}_{j}\leftarrow P^{t}_{j}\cup v$ (14)
-Reward: Our aim is to improve the value of the overall partitioning objective.
+3.3 Reward, Training & Inference
+Reward.
+Our aim is to improve the value of the overall partitioning objective.
 One way is to define the reward $R^{t}$ at step $t\geq 0$ as the change in objective value of the partitioning i.e $Obj(\mathcal{G},\mathcal{P}^{t})$ at step $t$, i.e., $R^{t}=(Obj(\mathcal{G},\mathcal{P}^{t+1})-Obj(\mathcal{G},\mathcal{P}^{t}))/(Obj(\mathcal{G},\mathcal{P}^{t+1})+Obj(\mathcal{G},\mathcal{P}^{t}))$.
@@
 However, this definition of reward focuses on short-term improvements instead of long-term.
-Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term.
+Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term (sutton2018reinforcement).
 The discounted rewards are computed as the sum of the rewards over a horizon of actions with varying degrees of importance (short-term and long-term).
 Mathematically,
-$D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}\vspace{-0.1in}$ (14)
+$D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}$ (15)
 where $T$ is the length of the horizon and $\gamma\in(0,1]$ is a discounting factor (hyper-parameter) describing how much we favor immediate rewards over the long-term future rewards.
 The above reward mechanism provides flexibility to our framework to be versatile to objectives of different nature, that may or may not be differentiable.
-This is an advantage over existing neural method (nazi2019gap) where having a differentiable form of the partitioning objective is a pre-requisite.
-3.2 Policy Loss Computation and Parameter Update
+This is an advantage over existing neural methods (nazi2019gap; tsitsulin2023graph) where having a differentiable form of the partitioning objective is a pre-requisite.
+Policy Loss Computation and Parameter Update.
 Our objective is to learn parameters of our policy network in such a way that actions that lead to an overall improvement of the partitioning objective are favored more over others.
@@
 ${\nabla J(\pi_{\theta})=\left[\sum_{t=0}^{T}\left(D^{t}\right)\nabla_{\theta}log\pi_{\theta}({a}^{t}/S^{t})\right]}$ (17)
-3.3 Training and Inference
-For a given set of training graphs, we optimize the parameters of the policy network $\pi_{\theta}$ for $T$ steps.
-Subsequently, given a new graph and partition objective, we adapt the policy parameters $\pi_{\theta}$ using Eq.
-16.
+Training and Inference.
+For a given graph, we optimize the parameters of the policy network $\pi_{\theta}$ for $T$ steps.
 Note that the trajectory length $T$ is not kept very large to avoid the long-horizon problem.
@@
 $\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{E}|+k)\times T^{\prime}\big{)}$.
-Here $\alpha$ is number of anchor nodes, $\beta$ is number of random walk iterations, $k$ is the number of partitions and $T^{\prime}$ is the number of iterations during inference.
+Here $\alpha$ is the number of anchor nodes, $\beta$ is number of random walk iterations, $k$ is the number of partitions and $T^{\prime}$ is the number of iterations during inference.
 Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$.
@@
 Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.
-In comparison to spectral clustering, which is a common graph partitioning algorithm having time complexity of $\mathcal{O}(|\mathcal{V}|^{3})$, the proposed technique’s complexity is much lower.
-This can also be attributed to the fact that NeuroCUT exploits the sparse nature of graphs unlike spectral clustering that performs matrix operations of $\mathcal{O}(|\mathcal{V}|^{3})$.
-For detailed derivation please see appendix  C.
-[…]
-[…]
-[…]
-Figure 2: Results of all methods on five datasets through radar plots.
-The color shade represents identity of each method.
-The legend in the first subplot is common for all plots.
-All values (for each partitioning objective) are normalized by dividing by the worst (maximum) cut value obtained on that dataset.
-The smaller the value, i.e, closer to the center, better is the performance.
-Further, when two polygons are overlapping, but not fully contained within each other, in general, the smaller polygon tends to indicate a better holistic performance across all objectives.
-Fig (a) represents cut values averaged across five datasets.
-Fig (b-f) depict results on each of the five datasets.
-The neural state-of-the-art GAP method produces the most inferior performance among considered baselines.
-Further in several cases it generates ”NaN” values (error).
-Our method NeuroCUT  either outperforms the baselines or produces the second best results while being close to the best one in all cases.
-Moreover, NeuroCUT  is the most robust and does not produce the worst cut values.
-The neural methods are trained and tested with the same number partitions ($k=5$).
-Unnormalized results for each dataset in table format are presented in Table  7 in appendix.
+For detailed derivation please see appendix  A.4.
 4 Experiments
 In this section, we demonstrate the efficacy of NeuroCUT against state-of-the-art methods and establish that:
-  • Efficacy and robustness: NeuroCUT consistently produces the best or the second best results in terms of the average performance over all objective functions.
-This establishes the robustness of NeuroCUT.
-  • Inductivity: As one of the major strengths, unlike the state-of-the-art neural model GAP (nazi2019gap), NeuroCUT is inductive on the number of partitions and, consequently, generalizes well over unseen number of partitions.
+  • Efficacy and robustness: NeuroCUT produces the best results over diverse partitioning objective functions. This establishes the robustness of NeuroCUT.
+  • Inductivity: As one of the major strengths, unlike existing neural models such as GAP (nazi2019gap), DMon (tsitsulin2023graph) and MinCutPool (bianchi2020spectral), our proposed method NeuroCUT is inductive on the number of partitions and consequently can generalize to unseen number of partitions.
 4.1 Experimental Setup
 Datasets:
-We use five datasets for our experiments among which one is synthetic. They are as follows:
-  • Cora and Citeseer (sen2008collective): These are citation networks.
-In these, a $k$-way cut could represent the partitioning of publications into $k$ clusters while minimizing the citations across the clusters.
-Each cluster might correspond to a distinct research field, shedding light on interdisciplinary connections or isolating specialized domains.
-  • Facebook (leskovec2012learning): In social network analysis, a $k$-way cut can delineate communities of users with minimal inter-community connections.
-This partitioning unveils hidden social circles and provides insights into how different groups interact within the larger network.
-The datasets have multiple connected components, we work with the largest connected component to make sure we can evaluate proper cuts.
-  • Harbin (li2019learning): In this road network, a $k$-way cut could entail partitioning road segments into $k$ subsets to optimize traffic flow between them.
-This partitioning aids in urban planning by identifying key intersections and routes that minimize congestion.
-  • Stochastic Block Models (SBM) (abbe2017community): This synthetic network is generated by SBM where $k$-way cuts reveal the underlying community structure.
-This helps to validate our method by comparing the identified partitions with the ground truth communities.
+We use four real datasets for our experiments. They are described below and their statistics are present in the appendix (please see Table 7).
+  • Cora and Citeseer (sen2008collective): These are citation networks where nodes correspond to individual papers and edges represent citations between papers.
+The node features are extracted using a bag-of-words approach applied to paper abstracts.
+  • Harbin (li2019learning): This is a road network extracted from Harbin city, China.
+The nodes correspond to road intersections and node features represent latitude and longitude of a road intersection.
+  • Actor (platonov2023critical): This dataset is based upon Wikipedia data where each node in the graph corresponds to an actor, and the edge between two nodes denotes co-occurrence on the same Wikipedia page.
+Node features correspond to keywords in Wikipedia pages.
+Partitioning objectives:
+We evaluate our method on a diverse set of four partitioning objectives described in Section  2, namely Normalized Cut, Balanced Cut, k-MinCut, and Sparsest Cut.
+In addition to evaluating on diverse objectives, we also choose a diverse number of partitions for evaluation, specifically, $k=2,\;5\;\text{and}\;10$.
 Baselines:
-We compare our proposed method NeuroCUT with the state-of-the-art neural method called GAP (nazi2019gap).
-It allows balanced normalized cuts with an end-to-end framework with a differentiable loss function which is a continuous relaxation version of normalized cut.
-We also use two popular non-neural baselines: hMetis (karypis1997multilevel) and spectral clustering (ng2001spectral).
+We compare our proposed method with both neural as well as non-neural methods.
+Neural baselines: We compare with DMon (tsitsulin2023graph), GAP (nazi2019gap), MinCutPool  (bianchi2020spectral) and Ortho (tsitsulin2023graph).
+DMon is the state-of-the-art neural attributed-graph clustering method.
+GAP is optimized for balanced normalized cuts with an end-to-end framework with a differentiable loss function which is a continuous relaxation version of normalized cut.
+MinCutPool optimizes for normalized cut and uses an additional orthogonality regularizer.
+Ortho is the orthogonality regularizer described in DMon and MinCutPool.
+Non-neural baselines: Following the settings of DMon (tsitsulin2023graph), we compare with K-means clustering applied on raw node features.
 Other settings:
 We run all our experiments on an Ubuntu 20.04 system running on Intel Xeon 6248 processor with 96 cores and 1 NVIDIA A100 GPU with 40GB memory for our experiments.
-We used GraphSage(hamilton2017inductive) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$.
+For NeuroCUT we used GraphSage(hamilton2017inductive) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$.
 We used Adam optimizer for training the parameters of our policy network $\pi_{\theta}$.
@@
 The default number of anchor nodes for computing positional embeddings is set to $35$.
+Dataset Method $k=2$ $k=5$ $k=10$
+Cora Kmeans 0.65 3.26 7.44
+MinCutPool 0.12 0.61 1.65
+DMon 0.57 3.07 7.40
+Ortho 0.80 1.88 4.06
+GAP 0.10 0.68 -
+NeuroCUT 0.02 0.33 0.92
+CiteSeer Kmeans 0.30 2.35 5.21
+MinCutPool 0.10 0.38 1.04
+DMon 0.33 2.71 6.84
+Ortho 0.28 1.51 3.25
+GAP 0.12 - -
+NeuroCUT 0.02 0.20 0.44
+Harbin Kmeans 0.56 2.34 5.40
+MinCutPool - - -
+DMon 0.98 3.25 -
+Ortho - - -
+GAP 0.25 - -
+NeuroCUT 0.01 0.07 0.28
+Actor Kmeans 0.99 4.00 8.98
+MinCutPool 0.55 1.97 4.73
+DMon 0.77 3.46 8.08
+Ortho 1.05 3.98 8.90
+GAP 0.20 - -
+NeuroCUT 0.17 0.99 4.66
+Table 2: Results on Normalized Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$.
+Dataset Method $k=2$ $k=5$ $k=10$
+Cora Kmeans 3.41 14.90 32.52
+MinCutPool 0.52 2.44 6.68
+DMon 0.45 1.89 5.82
+Ortho 2.73 7.06 15.70
+GAP 0.41 2.80 -
+NeuroCUT 0.13 1.46 3.03
+CiteSeer Kmeans 1.53 13.70 22.20
+MinCutPool 0.31 1.32 3.62
+DMon 0.35 1.21 3.62
+Ortho 0.88 4.15 10.60
+GAP 0.60 - -
+NeuroCUT 0.11 0.49 1.19
+Harbin Kmeans 1.91 7.44 17.03
+MinCutPool - - -
+DMon 2.01 - -
+Ortho - - -
+GAP 1.56 - -
+NeuroCUT 0.06 0.23 0.82
+Actor Kmeans 5.43 19.40 40.34
+MinCutPool 2.44 8.51 19.81
+DMon 1.71 9.50 21.01
+Ortho 3.42 16.55 40.4
+GAP 1.35 - -
+NeuroCUT 0.65 2.04 2.88
+Table 3: Results on Sparsest Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
 4.2 Results on Transductive Setting
-We compare our proposed method NeuroCUT against all the baselines where the neural models are trained and tested with the same number of partitions ($k=5$).
-Fig.
-2 presents the results for all datasets and all methods.
-The results are depicted using spider plots where the smaller the value i.e closer to the center, better is the performance.
-Further, in the plots, if two polygons are overlapping, but not fully contained within each other, in general the smaller polygon tends to indicate a better holistic performance across all objectives (heckert2002handbook).
-NeuroCUT outperforms the baselines or produces the next to optimal results.
-Moreover, NeuroCUT is the most robust among all the methods and does not produce the worst cut values in any of the settings.
-Overall, the performance of NeuroCUT is the most consistent among all the methods.
-Unnormalized results for each dataset in table format are presented in Table  7 in appendix.
-NeuroCUT incorporates causality during inference and this leads to the robust performance observed in fig.
-2.
-Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in GAP (nazi2019gap).
-We observe that the non-neural methods fail to generalize well across all objective functions since they do not learn the base heuristic as a function of the objective and the dataset.
-Finally, we note that in SBMs the task is easier as there are already $5$ clusters which aligns well with the the desired number of partitions, which is also $5$.
-4.3 Results on Inductive Setting
-Dataset Method $normalised$ $sparsest$ $balanced$ $kmin$
-Cora hMetis 0.772 0.772 3.177 0.097
-Spectral 0.723 1.503 1.845 0.008
-NeuroCUT 0.591 2.451 0.981 0.022
-CiteSeer hMetis 0.415 0.415 1.554 0.056
-Spectral 0.328 1.094 0.792 0.003
-NeuroCUT 0.363 1.060 0.457 0.031
-SBMs hMetis 3.259 3.260 65.955 0.393
-Spectral 2.495 2.513 50.805 0.242
-NeuroCUT 3.241 5.505 45.602 0.038
-Table 2: Inductive results based on the number of partitions of all methods on all six datasets.
-The best ones are shown by green and the second best ones are in gray.
-Lower values are better.
-Our method NeuroCUT either outperforms the baselines or produces the second best results while being close to the best one in ten out of twelve settings.
-NeuroCUT is trained on different partition sizes ($k=5$ and $k=10$) and the test data involves an unseen partition size ($k=8$).
-GAP is not inductive in terms of number of partitions and thus the results are not present here.
-We show that NeuroCUT generalizes to unknown number of partitions.
-We compare it against only the non-neural baselines since GAP does not generalize to unseen number of partitions.
-Here, NeuroCUT is trained on different partition sizes ($k=5$ and $k=10$) and the test data involves an unseen partition size ($k=8$).
-Table 2 shows the results.
-Our method NeuroCUT either outperforms the baselines or produces the second best results while being close to the best one in most of the settings.
-Note that while the non-neural methods need to be re-run for the unseen partitions, NeuroCUT only runs forward pass to produce the results on an unseen partition size.
-4.4 Generalization to Topological changes in graphs
-In this section, we analyze the performance of NeuroCUT to topological changes in the graph.
-Specifically, once a model is trained, we aim to evaluate its performance when the topology of the graph changes which is a common scenario in dynamic networks such as social networks where connections are formed/deleted over time.
-Towards this first we train NeuroCUT on a given graph.
-Next, we add/remove $x\%$ percentage of edges in the graph.
-Then we predict the cuts on the perturbed graph.
-To predict cuts for the perturbed graph, our model NeuroCUT only needs to just run a forward pass over it.
-Tables  3,4,5,6 show the results on two datasets, namely, Cora and Facebook for different partitioning objectives at different degrees of perturbations.
-In almost all settings, our method NeuroCUT outperforms the baselines.
-The only neural baseline i.e GAP had to be retrained on perturbed graph and still produced NaN values(indicated by -) in most of the cases.
-These results show that our proposed method NeuroCUT is more robust in comparison to existing neural methods at different levels of perturbations.
-Additionally, in several cases it outperforms non-neural methods which were directly optimized on perturbed graphs.
-[…]
-Cora hMetis No 0.351 0.355 0.399 0.435 0.506
-Spectral No 0.346 0.442 0.479 0.424 0.490
-GAP Yes 0.512 0.500 0.483 0.808 0.664
-NeuroCUT Yes 0.321 0.307 0.364 0.428 0.569
-Facebook hMetis No 1.112 1.113 1.126 1.126 1.131
-Spectral No 2.034 1.676 1.918 1.92 1.679
-GAP Yes - - 0.482 - -
-NeuroCUT Yes 0.487 0.459 0.463 0.494 0.864
-Table 3: Perturbing $x\%$ Edges - Normalised Cuts
-[…]
-Cora hMetis No 1.442 1.464 1.649 1.819 2.169
-Spectral No 0.795 1.058 1.224 0.957 1.452
-GAP Yes 2.029 1.978 1.915 3.167 2.816
-NeuroCUT Yes 1.015 1.317 1.479 1.546 2.053
-Facebook hMetis No 51.669 52.937 52.110 52.962 52.236
-Spectral No 5.625 4.722 5.591 5.591 4.867
-GAP Yes - - 11.886 - -
-NeuroCUT Yes 3.528 3.516 3.745 4.060 5.298
-Table 4: Perturbing $x\%$ Edges - Sparsest Cuts
-[…]
-Cora hMetis No 0.352 0.355 0.399 0.435 0.506
-Spectral No 1.089 1.223 0.800 1.207 0.841
-GAP Yes 0.536 0.529 0.515 0.870 0.681
-NeuroCUT Yes 0.526 0.544 0.613 0.651 0.836
-Facebook hMetis No 1.112 1.113 1.126 1.127 1.131
-Spectral No 2.818 2.459 2.697 2.697 2.456
-GAP Yes - - 0.506 - -
-NeuroCUT Yes 0.831 0.793 0.800 0.855 0.921
-Table 5: Perturbing $x\%$ Edges - Balanced Cuts
-[…]
-Cora hMetis No 0.071 0.072 0.080 0.087 0.101
-Spectral No 0.002 0.001 0.054 0.001 0.064
-GAP Yes 0.099 0.097 0.094 0.114 0.126
-NeuroCUT Yes 0.010 0.013 0.014 0.018 0.021
-Facebook hMetis No 0.199 0.203 0.201 0.203 0.201
-Spectral No 0.000 0.000 0.000 0.000 0.000
-GAP Yes 0.024 0.026 0.044 0.026 0.027
-NeuroCUT Yes 0.000 0.000 0.000 0.000 0.000
-Table 6: Perturbing $x\%$ Edges - KMin Cuts
+In the transductive setting, we compare our proposed method NeuroCUT against the mentioned baselines, where the neural models are trained and tested with the same number of partitions.
+Tables 2-5 present the results for different partitioning objectives across all datasets and methods.
+For the objectives under consideration, a smaller value depicts better performance.
+NeuroCUT demonstrates superior performance compared to both neural and non-neural baselines across four distinct partitioning objectives, highlighting its robustness.
+Further, we would also like to point out that in many cases, existing baselines produce “nan” values which are represented by “-” in the result tables.
+Our method NeuroCUT incorporates dependency during inference and this leads to robust performance.
+Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in methods such as GAP (nazi2019gap).
+Further, unlike other methods, NeuroCUT also incorporates positional information in the form of Lipschitz embedding to better contextualize global node positional information.
+We also observe that the non-neural method K-means fails to perform on all objective functions since its objective is not aligned with the main objectives under consideration.
+Dataset Method $k=2$ $k=5$ $k=10$
+Cora Kmeans 0.32 0.34 0.61
+MinCutPool 0.06 0.12 0.17
+DMon 0.05 0.09 0.14
+Ortho 0.33 0.34 0.38
+GAP 0.05 0.10 0.11
+NeuroCUT $\sim$ 0.0 $\sim$ 0.0 0.06
+CiteSeer Kmeans 0.12 0.27 0.43
+MinCutPool 0.04 0.08 0.10
+DMon 0.05 0.07 0.12
+Ortho 0.13 0.47 0.30
+GAP 0.05 0.08 0.10
+NeuroCUT 0.01 0.02 0.03
+Harbin Kmeans 0.28 0.46 0.54
+MinCutPool 0.0 0.0 0.0
+DMon 0.10 0.01 $\sim$ 0.0
+Ortho 0.0 0.0 0.0
+GAP 0.01 0.01 0.01
+NeuroCUT $\sim$ 0.0 0.01 0.03
+Actor Kmeans 0.48 0.54 0.80
+MinCutPool 0.25 0.35 0.42
+DMon 0.17 0.39 0.44
+Ortho 0.35 0.65 0.83
+GAP 0.09 0.12 0.26
+NeuroCUT $\sim$ 0.0 $\sim$ 0.0 $\sim$ 0.0
+Table 4: Results on k-MinCut.
+In most of the cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$.
+Values less than $10^{-2}$ are approximated to 0.
+Dataset Method $k=2$ $k=5$ $k=10$
+Cora Kmeans 0.68 3.90 7.44
+MinCutPool 0.13 0.60 1.62
+DMon 0.11 0.48 1.47
+Ortho 0.80 1.89 4.06
+GAP 0.10 0.74 -
+NeuroCUT 0.45 0.64 1.08
+CiteSeer Kmeans 0.42 2.64 5.30
+MinCutPool 0.09 0.38 1.04
+DMon 0.10 0.37 1.1
+Ortho 0.28 1.51 3.32
+GAP 0.23 - -
+NeuroCUT 0.07 0.24 0.60
+Harbin Kmeans 0.56 2.37 5.41
+MinCutPool - - -
+DMon 1.30 - -
+Ortho - - -
+GAP 0.72 - -
+NeuroCUT 0.21 0.11 0.27
+Actor Kmeans 1.00 4.06 9.08
+MinCutPool 0.55 1.96 4.71
+DMon 0.34 2.01 4.80
+Ortho 1.05 3.90 8.91
+GAP 0.25 - -
+NeuroCUT 0.59 1.69 4.42
+Table 5: Results on Balanced Cut. In most cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
+4.3 Results on Inductivity to Partition Count
+Dataset Method Normalized Sparsest k-MinCut Balanced
+Cora Kmeans 7.44 32.52 0.61 7.53
+NeuroCUT-T 0.92 3.03 0.06 1.08
+NeuroCUT-I 1.18 5.04 0.02 2.03
+CiteSeer Kmeans 5.21 22.20 0.43 5.30
+NeuroCUT-T 0.44 1.19 0.03 0.60
+NeuroCUT-I 0.63 1.62 0.03 0.62
+Harbin Kmeans 5.40 17.03 0.54 5.41
+NeuroCUT-T 0.28 0.82 0.03 0.27
+NeuroCUT-I 0.27 0.87 0.03 0.29
+Table 6: Inductivity to unseen partition count.
+Here we set the target number of partitions $k=10$.
+I stands for inductive and T for transductive setting.
+In this table, the transductive version of NeuroCUT trained on $k=10$ serves as a point of reference when assessing the quality of the inductive version of NeuroCUT.
 […]
 […]
-Figure 3: Results on the initial and the final cut values at $k=5$.
+Figure 2: Results on the initial warm start and the final cut values obtained by NeuroCUT at $k{=}10$.
 It shows that our neural model NeuroCUT (Final) performs more accurate node and partition selection to optimize the objective function.
 Subsequently, there is a significant difference between the initial and final cut values.
+As detailed in Section 3.2, the decoupling of parameter size and the number of partitions allows NeuroCUT to generalize effectively to an unseen number of partitions.
+In this section, we empirically analyze the generalization performance of NeuroCUT to an unknown number of partitions.
+We compare it with the non-neural baseline K-means, as the neural methods like GAP, DMon, Ortho, and MinCutPool cannot be employed to infer on an unseen number of partitions.
+In Table 6, we present the results on inductivity.
+Specifically, we trained NeuroCUT on different partition sizes ($k=5$ and $k=8$) and tested it on an unseen partition size $k=10$.
+The inductive version, referred to as NeuroCUT-I [^2] For clarity purposes, in this experiment we use NeuroCUT-T to refer to the transductive setting whose results are presented in Sec.
+4.2, obtains high-quality results on different datasets.
+Note that while the non-neural method such as K-means needs to be re-run for the unseen partitions, NeuroCUT only needs to perform forward pass to produce the results on an unseen partition size.
 4.4 Ablation Studies
-Initial vs Final Cut Values.
-NeuroCUT takes a warm start by partitioning nodes based on clustering over their Lipschitz embeddings, which are subsequently fine-tuned auto-regressively through the reinforcement learning based framework.
-How much does the partitioning objective improve since the initial clustering?
+Initial Warm Start vs Final Cut Values.
+NeuroCUT takes a warm start by partitioning nodes based on clustering over their raw node features and Lipschitz embeddings, which are subsequently fine-tuned auto-regressively through the phase 1 and phase 2 of NeuroCUT.
+In this section, we measure, how much the partitioning objective has improved since the initial clustering?
 Figure 2 sheds light on this question.
-Specifically, it shows the difference between initial cut value due after clustering and the final cut value from the partitions produced by our method.
+Specifically, it shows the difference between the initial cut value after clustering and the final cut value from the partitions produced by our method.
 We note that there is a significant difference between the initial and final cut values, which essentially shows the effectiveness of NeuroCUT.
+Figure 3: Node Selection in Phase 1 with Heuristic vs Random: Relative % improvement (gain) in cut values obtained by NeuroCUT when using node selection heuristic against random node selection for $k=5$.
+In most cases, our heuristic finds significantly better cuts than a random node selection procedure.
 […]
-[…]
-Figure 4: NeuroCUT results with heuristics for node selection in Phase 1 for $k=5$.
-Our heuristic finds better cuts than a random selection algorithm for two objective functions: Normalized Cut and KMin Cut.
-Impact of Node Selection Procedures.
-Figure 4 shows how our proposed heuristic for node selection helps in finding the near-optimal cuts for any given objective function.
-We compare and show that a simpler node selection where we select all the nodes one by one in a random order produces substantially inferior results.
-Furthermore, the difference lies not only in final cut-values but also we require much less perturbations.
-In App. D of appendix, we present additional results on variation of parameters of NeuroCUT.
 5 Conclusion
-In this paper, we have introduced a novel graph reinforcement learning framework NeuroCUT to effectively address the graph partitioning problem.
-A distinctive feature of our framework is its adaptability to diverse partitioning objectives that may or may not be differentiable, which sets it apart from conventional algorithms.
-Our approach has also shown the ability to generalize on unseen graphs as well as unseen number of partitions.
-The efficacy of our approach is empirically validated through an extensive evaluation on six datasets with four graph partitioning objectives.
-Notably, our method shows remarkable performance when compared to both non-neural and neural state-of-the-art techniques, proving its competence in both inductive and transductive settings.
+In this work, we study the problem of graph partitioning with node features.
+Existing neural methods for addressing this problem require the target objective to be differentiable and necessitate prior knowledge of the number of partitions.
+In this paper, we introduced NeuroCUT, a framework to effectively address the graph partitioning problem with node features.
+NeuroCUT tackles these challenges using a reinforcement learning-based approach that can adapt to any target objective function.
+Further, attributed to its decoupled parameter space and partition count, NeuroCUT can generalize to an unseen number of partitions.
+The efficacy of our approach is empirically validated through an extensive evaluation on four datasets, four graph partitioning objectives and diverse partition counts.
+Notably, our method shows significant performance gains when compared to the state-of-the-art techniques, proving its competence in both inductive and transductive settings.
 In summary, NeuroCUT offers a versatile, inductive, and empirically validated solution for graph partitioning.
-Dataset $|\mathcal{V}|$ $|\mathcal{E}|$ Method $normalised$ $sparsest$ $balanced$ $kmin$ AVG DIFF MAX DIFF
-ACTDIFF ACTDIFF ACTDIFF ACTDIFF
-Cora 2485 5069 hMetis 0.343   0.115 1.412   0.061 0.345   0.000 0.069   0.063 0.060 0.115
-Spectral 0.370   0.142 1.351   0.000 1.165   0.820 0.006   0.000 0.240 0.820
-GAP 0.505   0.277 2.074   0.723 0.536   0.191 0.095   0.089 0.320 0.723
-NeuroCUT 0.228   0.000 1.351   0.000 0.570   0.225 0.011   0.005 0.057 0.225
-CiteSeer 2120 3679 hMetis 0.198   0.025 0.673   0.227 0.198   0.000 0.038   0.037 0.072 0.227
-Spectral 0.173   0.000 0.446   0.000 0.929   0.731 0.001   0.000 0.182 0.731
-GAP 0.267   0.094 0.795   0.349 0.312   0.114 0.043   0.042 0.150 0.349
-NeuroCUT 0.211   0.038 0.712   0.266 0.296   0.098 0.036   0.035 0.109 0.266
-SBMs 500 5150 hMetis 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
-Spectral 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
-GAP - - - 0.035   $0$
-NeuroCUT 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
-Facebook 1034 26749 hMetis 1.115   0.858 52.319   48.198 1.115   0.143 0.200   0.197 12.349 48.198
-Spectral 1.670   1.413 4.722   0.601 2.459   1.487 0.003   0.000 0.875 1.487
-GAP - - - 0.025   0.022 - -
-NeuroCUT 0.257   0.000 4.121   0.000 0.972   0.000 0.015   0.012 0.003 0.012
-Harbin 6235 15205 hMetis 0.058 0.000 0.184   0.000 0.058   0.000 0.011   0.000 0.000 0.000
-Spectral 0.104   0.046 0.348   0.164 0.403   0.345 0.015   0.004 0.140 0.345
-GAP 0.175   0.117 0.551   0.367 0.269   0.211 0.017   0.006 0.175 0.367
-NeuroCUT 0.092   0.034 0.286   0.102 0.099   0.041 0.018   0.007 0.046 0.102
-Table 7: Results of all methods on all five datasets.
-“ACT” denotes the actual values of the measures and “DIFF” shows the difference with the best value for the dataset.
-“AVG DIFF” and “MAX DIFF” present the average and the maximum of the differences over all the measures respectively.
-The best average differences are shown by green and the second best ones are in gray.
-The worst result, i.e., the maximum difference is shown in red.
-The neural state-of-the-art GAP method produces the most inferior performance among considered baselines.
-Our method NeuroCUT either outperforms the baselines (green) or produces the second best (gray) results while being close to the best one in all cases.
-Moreover, NeuroCUT is the most robust and never produces the worst maximum difference (i.e., in red).
-The neural methods are trained and tested with the same number partitions ($k=5$).
-Appendix A Result for all datasets in transductive setting
-In Table  7 we present the results obtained by all methods on different datasets. These correspond to radar plot obtained in fig. 2 in main paper .
-Appendix B Clustering for Initialization
+Appendix A Appendix
+A.1 Additional Related Work
+In recent years there has been a growing interest in using machine learning techniques to design heuristics for NP-Hard combinatorial problems on graphs.
+This line of work is motivated from the observation that traditional algorithmic approaches are limited to using the same heuristic irrespective of the underlying data distribution.
+Neural approaches, instead, focus on learning the heuristic as a function of the data without deriving explicit algorithms (duong2023deep; jung2022learning; nazi2019gap).
+Significant advances have been made in learning to solve a variety of combinatorial optimization problems such as Set Cover, Travelling Salesman Problem, Influence Maximization etc.
+(khalil2017learning; gcomb; kool2018attention; joshi2019efficient).
+The neural architectural designs in these methods are generally suitable for a specific set of CO problems (mazyavkina2021reinforcement) of similar nature.
+For instance, AM  (kool2018attention) and NeuroLKH (xin2021neurolkh) are targeted towards routing problems.
+GCOMB (gcomb) and LeNSE (ireland2022lense) aim to solve problems that are submodular in nature.
+Owing to these assumptions, the architectural designs in these works are not suitable for our problem of $k$-way inductive graph partitioning with versatile partitioning objectives.
+A.2 Clustering for Initialization
 As discussed in sec 3.1 in main paper, we first cluster the nodes of the graph into $k$ clusters where $k$ is the number of partitions.
-Towards this, we apply K-means algorithm (macqueen1967some) on the nodes of the graph where a node is represented by its positional node embedding i.e $pos(u)\;\forall u\in\mathcal{V}$.
+Towards this, we apply K-means algorithm (macqueen1967some) on the nodes of the graph where a node is represented by its raw node features and positional representation i.e $\mathbf{emb_{init}}(u)\;\forall u\in\mathcal{V}$ based upon eq.
+8.
 Further, we used $L_{\inf}$ norm as the distance metric for clustering.
-Appendix C Time complexity of NeuroCUT
+A.3 Dataset Statistics
+Table  7 describes the statistics of datasets used in our work. For all datasets we use their largest connected component.
+Dataset $|\mathcal{V}|$ $|\mathcal{E}|$ Number of features
+Cora 2485 5069 1433
+CiteSeer 2120 3679 3703
+Harbin 6598 10492 2
+Actor 6198 14879 931
+Table 7: Dataset Statistics
+A.4 Time complexity of NeuroCUT
 We derive the time complexity of forward pass of NeuroCUT as discussed in sec. 3.4 in main paper.
@@
 (3) The node selection algorithm is used that computes score for each node based upon its neighborhood using eq.
-9.
+10.
 This consumes $\mathcal{O}(|\mathcal{V}|\times d)$ time.
 (4) Finally for the selected node, its partition has to be determined using eq.
-11 and  12.
+12 and  13.
 This takes $\mathcal{O}(d+k)$ time, as we consider only the neighbors of the selected node to compute partition score.
@@
 Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.
-Appendix D Impact of number of anchor nodes
-In this section we analyze the impact of number of anchor nodes chosen for computing positional embeddings of nodes in the graph $pos(u)$.
-Fig.
-5 shows the results where $Y$-axis shows the normalized cut value and $X$-axis shows the number of anchor nodes.
-The best results are achieved at the number of anchor nodes being 35 except in Citeseer.
-The other variation in the number of anchor nodes produces similar results often.
-Figure 5: Results on varying the number of anchor nodes on the normalized cut value produced by NeuroCUT.

### v3

@@
-Written by AAAI Press Staff1
-Abstract
+NeuroCUT: A Neural Approach for Robust Graph Partitioning
+Abstract.
 Graph partitioning aims to divide a graph into $k$ disjoint subsets while optimizing a specific partitioning objective.
@@
 Through empirical evaluation, we demonstrate that NeuroCUT excels in identifying high-quality partitions, showcases strong generalization across a wide spectrum of partitioning objectives, and exhibits strong generalization to unseen partition count.
-Abstract
-1 Introduction and Related Work
-Graph partitioning is a fundamental problem in network analysis with numerous real-world applications in various domains such as system design in online social networks (pujol2009divide), dynamic ride-sharing in transportation systems (tafreshian2020trip), VLSI design (kahng2011vlsi), and preventing cascading failure in power grids (li2005strategic).
+Graph Partitioning, Min Cut, Robustness, Versatile Partitioning Objectives, Inductive learning, GNN, Reinforcement learning.
+copyright: acmcopyrightjournalyear: 2023doi: 10.1145/3637528.3671815journalyear: 2024copyright: acmlicensedconference: 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spainbooktitle: 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08doi: 10.1145/3637528.3671815ccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligenceccs: Mathematics of computing Graph algorithmsccs: Mathematics of computing Graph algorithms
+1. Introduction
+Graph partitioning is a fundamental problem in network analysis with numerous real-world applications in various domains such as system design in online social networks (Pujol et al., 2009), dynamic ride-sharing in transportation systems (Tafreshian and Masoud, 2020), VLSI design (Kahng et al., 2011), and preventing cascading failure in power grids (Li et al., 2005).
 The goal of graph partitioning is to divide a given graph into disjoint subsets where nodes within each subset exhibit strong internal connections while having limited connections with nodes in other subsets.
 Generally speaking, the aim is to find somewhat balanced partitions while minimizing the number of edges across partitions.
-Several graph partitioning formulations have been studied in the literature, mostly in the form of discrete optimization (karypis1997multilevel; karypis1998multilevelk; karypis1999multilevel; andersen2006local; chung2007four).
-The majority of the formulations are NP-hard and thus the proposed solutions are either heuristics or algorithms with approximate solutions (karypis1997multilevel).
+1.1. Related Work
+Several graph partitioning formulations have been studied in the literature, mostly in the form of discrete optimization (Karypis et al., 1997; Karypis and Kumar, 1998, 1999; Andersen et al., 2006; Chung, 2007).
+The majority of the formulations are NP-hard and thus the proposed solutions are either heuristics or algorithms with approximate solutions (Karypis et al., 1997).
+Among these, two widely used methods are Spectral Clustering (Ng et al., 2001) and hMETIS (Karypis et al., 1997).
+Spectral Clustering partitions a graph into clusters based on the eigenvectors of a similarity matrix derived from the graph.
+hMETIS is a hypergraph partitioning method that divides a graph into clusters by maximizing intra-cluster similarity while minimizing inter-cluster similarity.
+However, such techniques are confined to specific objective functions and are unable to leverage the available node features in the graph.
 Recently, there have been attempts to solve graph partitioning problem via neural approaches.
@@
 Node features supply additional information and provide contextual insights that may improve the accuracy of graph partitioning.
-For instance, (wang2017mgae) has recognized the importance of incorporating node attributes in graph clustering where nodes are partitioned into disjoint groups.
-Note that there are existing neural approaches (khalil2017learning; gcomb; kool2018attention; joshi2019efficient) to solve other NP-hard graph combinatorial problems (e.g., minimum vertex cover).
+For instance, (Wang et al., 2017) has recognized the importance of incorporating node attributes in graph clustering where nodes are partitioned into disjoint groups.
+Note that there are existing neural approaches (Khalil et al., 2017; Manchanda et al., 2020; Kool et al., 2018; Manchanda et al., 2022, 2024; Joshi et al., 2019; Tian et al., 2024; Gasse et al., 2019; Ranjan et al., 2022; Manchanda and Ranu, 2023; Goyal et al., 2020) to solve other NP-hard graph combinatorial problems (e.g., minimum vertex cover).
 However these methods are not generic enough to solve graph partitioning.
+For example, although S2VDQN (Khalil et al., 2017) learns to solve the Maxcut problem, it is designed for the specific case of bi-partitioning.
+This hinders its applicability to the target problem setup of $k$-way partitioning [^1] Additional related work is present in Appendix LABEL:app:additional_rel_work.
 In this paper, we build a single framework to solve several graph partitioning problems.
-One of the most relevant to our work is the method DMoN by muller2023graph.
+One of the most relevant to our work is the method DMoN by Tsitsulin et al.
+(2023a).
 This method designs a neural architecture for cluster assignments and use a modularity-based objective function for optimizing these assignments.
-Another method, that is relevant to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (nazi2019gap).
+Another method, that is relevant to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (Nazi et al., 2019).
 It proposes a differentiable loss function for partitioning based on a continuous relaxation of the normalized cut formulation.
-Deep-MinCut being an unsupervised approach learns both node embeddings and the community structures simultaneously where the objective is to minimize the mincut loss (duong2023deep).
-Another method solves the multicut problem where the number of partitions is not an input to the problem (jung2022learning).
+Deep-MinCut being an unsupervised approach learns both node embeddings and the community structures simultaneously where the objective is to minimize the mincut loss (Duong et al., 2023).
+Another method solves the multicut problem where the number of partitions is not an input to the problem (Jung and Keuper, 2022).
 The idea is to construct a reformulation of the multicut ILP constraints to a polynomial program as a loss function.
 However, the problem formulation is different than the generic normalized cut or mincut problem.
-Finally,  (gatti2022graph) solves the normalized cut problem only for the case where the number of partitions is exactly two.
+Another related work is DGCluster (Bhowmick et al., 2024), which proposes to solve the attributed graph clustering problem while maximizing modularity when $k$ is not known beforehand.
+Finally,  (Gatti et al., 2022) solves the normalized cut problem only for the case where the number of partitions is exactly two.
 Nevertheless, these neural approaches for the graph partitioning problem often do not use node features and only limited to a distinct partitioning objective.
@@
 Hence, it is important for a learned model to generalize to any partition count without retraining.
+In $k$-way graph partitioning, a model demonstrates inductivity to the number of partitions when it can infer on varying partition numbers without specific training for each.
 Existing neural approaches are non-inductive to the number of partitions, i.e they can only infer on number of partitions on which they are trained.
 Additionally, it is worth noting that the optimal number of partitions is often unknown beforehand.
+This capability is crucial in practical applications like chip design, where graph partitioning optimizes logic cell placement by dividing netlists (circuits) into smaller subgraphs, aiding independent placement.
+As the optimal partition count is frequently unknown in advance, experimenting with different partition numbers is a common practice.
 Hence, it is a common practice to experiment with different partition counts and evaluate their impact on the partitioning objective.
-While the existing methods (muller2023graph; nazi2019gap; gatti2022graph) assume that the number of partitions ($k$) is known beforehand, our proposed method can generalize to any $k$.
+While the existing methods (Tsitsulin et al., 2023a; Nazi et al., 2019; Gatti et al., 2022) assume that the number of partitions ($k$) is known beforehand, our proposed method can generalize to any $k$.
   • Non-generalizability to different cut functions: Multiple objective functions for graph cut have been studied in the partitioning literature.
 The optimal objective function hinges upon the subsequent application in question.
-For instance, the two most relevant studies, DMoN (muller2023graph) and GAP (nazi2019gap) focus on maximizing modularity and minimizing normalized cut respectively.
+For instance, the two most relevant studies, DMoN (Tsitsulin et al., 2023a) and GAP (Nazi et al., 2019) focus on maximizing modularity and minimizing normalized cut respectively.
 Our framework is generic and can solve different partitioning objectives.
@@
 As we illustrate in § 2, the assumption does not always hold in the real-world.
-For instance, the sparsest cut (chawla2006hardness) and balanced cut (nazi2019gap) objectives are not differentiable.
-1.1 Contributions:
+For instance, the sparsest cut (Chawla et al., 2006) and balanced cut (Nazi et al., 2019) objectives are not differentiable.
+1.2. Our Contributions
 In this paper, we circumvent the above-mentioned limitations through the following key contributions.
@@
 Our empirical investigation substantiates the efficacy of NeuroCUT in partitioning tasks, showcasing its robustness across a spectrum of objective functions.
-We also demonstrate the capacity of NeuroCUT to generalize effectively to partition sizes that it has not encountered during training.
-2 Problem Formulation
+We also demonstrate the capability of NeuroCUT to generalize effectively to partition sizes that have not been seen during training.
+2. Problem Formulation
 In this section, we introduce the concepts central to our work and formulate the problem. All the notations used in this work are outlined in Table 1.
-Definition 1 (Graph)
+Definition 1 (Graph).
 We denote a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E},\textbf{{X}})$ where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges and $\textbf{{X}}\in\mathbb{R}^{|\mathcal{V}|\times|F|}$ refers to node feature matrix where $F$ is the set of all features in graph $\mathcal{G}$.
-Definition 2 (Cut)
+Definition 2 (Cut).
 A cut $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is a partition of $\mathcal{V}$ into two subsets $\mathcal{S}$ and $\mathcal{T}$.
-The cut-set of $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is the set $(u,v)\in\mathcal{E}|u\in\mathcal{S},v\in\mathcal{T}$ of edges that have one endpoint in $\mathcal{S}$ and the other endpoint in $\mathcal{T}$.
-Definition 3 (Graph Partitioning)
+The cut-set of $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is the set ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{}(u,v)\in\mathcal{E}|u\in\mathcal{S},v\in\mathcal{T}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\}}$ of edges that have one endpoint in $\mathcal{S}$ and the other endpoint in $\mathcal{T}$.
+Definition 3 (Graph Partitioning).
 Given graph $\mathcal{G}$, we aim to partition $\mathcal{G}$ into $k$ disjoint sets $\mathcal{P}=\{P_{1},P_{2},\cdots,P_{k}\}$ such that the union of the nodes in those sets is equal to $\mathcal{V}$ i.e $\cup_{i=1}^{k}P_{i}=\mathcal{V}$ and each node belongs to exactly one partition.
@@
 [^2] Our framework is not restricted to these objectives.
-  1. 1.
-    $k$-MinCut (saran1995finding): Partition a graph into $k$ partitions such that the total number of edges across partitions is minimized.
+  1. (1)
+    $k$-MinCut (Saran and Vazirani, 1995): Partition a graph into $k$ partitions such that the total number of edges across partitions is minimized.
 […]
     Here $P_{l}$ refers to the set of elements in $l^{th}$ partition of $\mathcal{P}$ as described in Def.  3 and $\overline{P_{l}}$ refers to set of elements not in $P_{l}$.
-  2. 2.
-    Normalized Cut (shi2000normalized): The $k$-MinCut criteria favors cutting small sets of isolated nodes in the graph.
+  2. (2)
+    Normalized Cut (Shi and Malik, 2000): The $k$-MinCut criteria favors cutting small sets of isolated nodes in the graph.
 To avoid this unnatural bias for partitioning out small sets of points, normalised cut computes the cut cost as a fraction of the total edge connections to all the nodes in the graph.
-    $\operatorname{Ncut}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}$ (2)
+    (2) $\operatorname{Ncut}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}$
     Here, $\operatorname{vol}\left(P_{l},\mathcal{V}\right):=\sum_{v_{i}\in P_{l},v_{j}\in\mathcal{V}}e(v_{i},v_{j})$.
-  3. 3.
-    Balanced Cuts (nazi2019gap): Balanced cut favours partitions of equal sizes so an extra term that indicates the squared distance from equal sized partition is added to normalised cuts.
-    $\operatorname{Balanced-Cuts}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}+\frac{(|P_{l}|-|\mathcal{V}|/k)^{2}}{|\mathcal{V}|^{2}}$ (3)
+  3. (3)
+    Balanced Cut (Nazi et al., 2019): Balanced cut favours partitions of equal sizes so an extra term that indicates the squared distance from equal sized partition is added to normalized cuts.
+    (3) $\operatorname{Balanced-Cuts}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}+\frac{(|P_{l}|-|\mathcal{V}|/k)^{2}}{|\mathcal{V}|^{2}}$
     Here, $\operatorname{vol}\left(P_{l},\mathcal{V}\right):=\sum_{v_{i}\in P_{l},v_{j}\in\mathcal{V}}e(v_{i},v_{j})$
-  4. 4.
-    Sparsest Cuts (chawla2006hardness): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition.
+  4. (4)
+    Sparsest Cut (Chawla et al., 2006): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition.
 We generalize it to $k$-way sparsest cuts by summing up the value for all the partitions.
 The intuition behind sparsest cuts is that any partition should neither be very large nor very small.
-    $\phi(S,\bar{S})$ $=\frac{\operatorname{cut}(S,\bar{S})}{\min(|S|,|\bar{S}|)}$ (4)
-    $\operatorname{Sparsest-Cuts}(\mathcal{\mathcal{P}})$ $=\sum_{l=1}^{|\mathcal{P}|}\phi(P_{l},\overline{P_{l}})$ (5)
-Problem 1 (Learning to Partition Graph )
+    (4) $\operatorname{Sparsest-Cuts}(\mathcal{\mathcal{P}})$ $=\sum_{l=1}^{|\mathcal{P}|}\phi(P_{l},\overline{P_{l}})$
+    (5) $where\;\;\phi(S,\bar{S})$ $=\frac{\operatorname{cut}(S,\bar{S})}{\min(|S|,|\bar{S}|)}$
+Problem 1 (Learning to Partition Graph ).
 Given a graph $\mathcal{G}$ and the number of partitions $k$, the goal is to find a partitioning $\mathcal{P}$ of the graph $\mathcal{G}$ that optimizes a target objective function ${Obj}(\mathcal{G},\mathcal{P})$.
@@
 In addition to our primary goal of finding a partitioning that optimizes a certain objective function, we also desire policy $\pi$ to have the following properties:
-  1. 1.
+  1. (1)
     Inductive: Policy $\pi$ is inductive if the parameters of the policy are independent of both the size of the graph and the number of partitions $k$.
 If the policy is not inductive then it will be unable to infer on unseen size graphs/number of partitions.
-  2. 2.
+  2. (2)
     Learning Versatile Objectives: To optimize the parameters of the policy, a target objective function is required.
@@
 Hence, the policy $\pi$ should be capable of learning to optimize for a target objective that may or may not be differentiable.
-Figure 1: Architecture of NeuroCUT.
-First, the initial partitioning of the graph is performed based on node raw features and positional embeddings.
+Figure 1.
+Architecture of NeuroCUT.
+First, the initial partitioning of the graph is performed based on node features and positional embeddings.
 These embeddings are refined using GNN to infuse toplogical information from neighborhood.
 At each step a node is selected and its partitioned is updated.
-During the training phase the GNN parameters are updated and hence embeddings are re-computed.
+During training the GNN parameters are updated and hence embeddings are re-computed.
 During inference, the GNN is called only once to compute the embeddings of the nodes of the graph.
+Table 1. Notations used in the paper
 Symbol Meaning
-$\mathcal{G}$
-Graph
-$\mathcal{V}$
-Node set
-$e$
-Edge $e\in\mathcal{E}$
-$\mathcal{E}$
-Edge set
-X
-Feature matrix containing raw node features
-$\mathcal{N}_{v}$
-Neighboring nodes of node $v$
-$Obj(\mathcal{G},\mathcal{P})$
-Objective function based upon graph $\mathcal{G}$ and its partitioning $\mathcal{P}$
-$k$
-Number of partitions
-$\mathcal{P}^{t}$
-Partitioning at time $t$
-${P}^{t}_{i}$
-$i^{th}$ partition at time $t$
-$\overline{{P}_{i}}$
-Set of nodes that are not in the $i^{th}$ partition
-$\mathcal{S}^{t}$
-State of system at step $t$
-$\operatorname{PART}(\mathcal{P}^{t},v)$
-Partition of node $v$ at time $t$
-$\mathcal{S}^{t}$
-State representation of Partitions and Graph at step $t$
-$\mathbf{pos}(v)$
-Positional embedding of node $v$
-$\mathbf{emb_{init}}(v)$
-Initial embedding of node $v$
-$\alpha$
-Number of anchor nodes for lipschitz embedding
-$T$
-Length of trajectory
-$\pi$
-Policy function
-Table 1: Notations used in the paper
-3 NeuroCUT: Proposed Methodology
+$\mathcal{G}$ Graph
+$\mathcal{V}$ Node set
+$e$ Edge $e\in\mathcal{E}$
+$\mathcal{E}$ Edge set
+X Feature matrix containing raw node features
+$\mathcal{N}_{v}$ Neighboring nodes of node $v$
+$Obj(\mathcal{G},\mathcal{P})$ Objective function based upon graph $\mathcal{G}$ and its partitioning $\mathcal{P}$
+$k$ Number of partitions
+$\mathcal{P}^{t}$ Partitioning at time $t$
+${P}^{t}_{i}$ $i^{th}$ partition at time $t$
+$\overline{{P}_{i}}$ Set of nodes that are not in the $i^{th}$ partition
+$\mathcal{S}^{t}$ State of system at step $t$
+$\operatorname{PART}(\mathcal{P}^{t},v)$ Partition of node $v$ at time $t$
+$\mathcal{S}^{t}$ State representation of Partitions and Graph at step $t$
+$\mathbf{pos}(v)$ Positional embedding of node $v$
+$\mathbf{emb_{init}}(v)$ Initial embedding of node $v$
+$\alpha$ Number of anchor nodes for lipschitz embedding
+$T$ Length of trajectory
+$\pi$ Policy function
+3. NeuroCUT: Proposed Methodology
 Fig.
@@
 After a node’s partition is updated, the reward with respect to change in partitioning objective value is computed and the parameters of the policy are optimized.
-The reward is learned through reinforcement learning (RL) (sutton2018reinforcement).
+The reward is learned through reinforcement learning (RL) (Sutton and Barto, 2018).
 The choice of using RL is motivated through two observations.
@@
 We next formalize each of these notions in our MDP formulation.
-3.1 State: Initialization & Positional Encoding
+3.1. State: Initialization & Positional Encoding
 Initialization.
@@
 The graph is clustered into $k$ clusters based on their raw features and positional embeddings (discussed below).
-We use K-means (macqueen1967some) algorithm for this task.
+We use K-means (MacQueen et al., 1967) algorithm for this task.
 The details of the clustering are present in Appendix A.1.
@@
 Two nodes that are closer in the graph, should be closer in the embedding space.
-Towards this, we use Lipschitz Embedding (bourgain1985lipschitz).
+Towards this, we use Lipschitz Embedding (Bourgain, 1985).
 Let $\mathcal{A}=\{a_{1},\cdots,a_{\alpha}\}\subseteq\mathcal{V}$ be a randomly selected subset of $\alpha$ nodes.
 We call them anchor nodes.
-From each anchor node $i$, the walker starts a random walk (bianchini2005inside) and jumps to a neighboring node $j$ with a transition probability ($\mathbf{W}_{ij}$) governed by the transition probability matrix $\mathbf{W}\in\mathbb{R}^{|V|\times|V|}$.
+From each anchor node $i$, the walker starts a random walk (Bianchini et al., 2005) and jumps to a neighboring node $j$ with a transition probability ($\mathbf{W}_{ij}$) governed by the transition probability matrix $\mathbf{W}\in\mathbb{R}^{|V|\times|V|}$.
 Furthermore, at each step, with probability $c$ the walker jumps to a neighboring node $j$ and returns to the node $i$ with $1-c$.
 Let $\vec{r}_{ij}$ corresponds to the probability of the random walker starting from node $i$ and reaching node $j$.
-$\vec{r}_{i}=c\tilde{\mathbf{W}}\vec{r}_{i}+(1-c)\vec{e}_{i}$ (6)
+(6) $\vec{r}_{i}=c\tilde{\mathbf{W}}\vec{r}_{i}+(1-c)\vec{e}_{i}$
 Eq.
@@
 Based upon the obtained random walk vectors for the set of anchor nodes $\mathcal{A}$, we embed all nodes $u\in\mathcal{V}$ in a $\alpha$-dimensional feature space:
-$\mathbf{pos}(u)=[r_{1u},r_{2u},\cdots,r_{\alpha u}]$ (7)
+(7) $\mathbf{pos}(u)=[r_{1u},r_{2u},\cdots,r_{\alpha u}]$
 To accommodate both raw feature and positional information we concatenate the raw node features i.e $\textbf{X}[u]$ with the positional embedding $\mathbf{pos}(u)$ for each node $u$ to obtain the initial embedding which will act as input to our neural model.
 Specifically,
-$\mathbf{emb_{init}}(u)=\textbf{X}[u]\;\;\mathbin{\|}\;\;\mathbf{pos}(u)$ (8)
+(8) $\mathbf{emb_{init}}(u)=\textbf{X}[u]\;\;\mathbin{\|}\;\;\mathbf{pos}(u)$
 […]
@@
 The state of the system at step $t$ is defined as
-$S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{emb_{init}}(v)\;\;\forall v\in P_{i}^{t}\}\}$ (9)
+(9) $S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{emb_{init}}(v)\;\;\forall v\in P_{i}^{t}\}\}$
 Here state of each partition is represented by the collection of initial embedding of nodes in that partition.
-3.2 Action: Selection of Nodes and Partitions
+3.2. Action: Selection of Nodes and Partitions
 Towards finding the best partitioning scheme for the target partition objective we propose a 2-step action strategy to update the partitions of nodes.
@@
 We discuss both these phases in details below.
-Phase 1: Node selection to identify node to perturb
+3.2.1. Phase 1: Node Selection to Identify Node to Perturb
 Let $\operatorname{PART}(\mathcal{P}^{t},v)$ denote the partition of the node $v$ at step $t$.
@@
 Specifically the score of node $v$ at step $t$ is defined as:
-$\operatorname{NodeScore}^{t}[v]=\\ {\frac{{\max_{p\in P^{t}-\operatorname{PART}(P^{t},v)}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=p|}}}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=\operatorname{PART}(P^{t},v)|}}$ (10)
-Further, a node having a higher degree implies it has several edges associated to it, hence, an incorrect placement of it could contribute to higher partitioning value.
-Hence, we normalize the scores by the $degree$ of the node.
-Phase 2: Inductive Method for Partition Selection
+(10) $\operatorname{NodeScore}^{t}[v]=\\ \frac{\max_{p\in{P}^{t}\setminus\operatorname{PART}({P}^{t},v)}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}({P}^{t},u)=p|}}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}({P}^{t},u)=\operatorname{PART}({P}^{t},v)|}\times\frac{1}{degree(v)}$
+For a node of interest $v$, the numerator computes the maximum number of neighbors in a different partition than that of $v$.
+As described in Table  1, $\operatorname{PART}(P^{t},v)$ refers to the partition of $v$ at step $t$.
+The expression $P^{t}\setminus\operatorname{PART}(P^{t},v)$ computes all other partitions except the partition of $v$.
+The term $|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=p|$ computes the number of neighbors of $v$ in the partition $p$.
+The denominator computes the number of neighbors of node $v$ in the same partition as $v$, referred to as $\operatorname{PART}(P^{t},v)$.
+Further, a node having a higher degree implies that it has several edges associated to it.
+Hence, an incorrect placement of it could contribute to a higher partitioning value.
+Therefore, we normalize the scores by the $degree$ of the node.
+3.2.2. Phase 2: Inductive Method for Partition Selection
 Once a node is selected, the next phase involves choosing the new partition for the node.
-Towards this, we design an approach empowered by Graph Neural Networks (GNNs) (hamilton2017inductive) which enables the model to be inductive with respect to size of graph.
-Further, instead of predicting a fixed-size score vector (nazi2019gap; tsitsulin2023graph) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too.
+Towards this, we design an approach empowered by Graph Neural Networks (GNNs) (Hamilton et al., 2017) which enables the model to be inductive with respect to size of graph.
+Further, instead of predicting a fixed-size score vector (Nazi et al., 2019; Tsitsulin et al., 2023b) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too.
 We discuss both above points in section below.
 Message Passing through Graph Neural Network
-To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (hamilton2017inductive).
+To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (Hamilton et al., 2017).
 GNNs combine node feature information and the graph structure to learn better representations via feature propagation and aggregation.
@@
 We perform $L$ layers of message passing to compute representations of nodes.
-To generate the embedding for node $u$ at layer $l+1$ we perform the following transformation(hamilton2017inductive):
-$\mathbf{h}_{u}^{l+1}$ $=\mathbf{W}^{l}_{1}\mathbf{h}_{u}^{l}+\mathbf{W}^{l}_{2}\cdot\frac{1}{|\mathcal{N}_{u}|}\sum_{u^{\prime}\in\mathcal{N}_{u}}\mathbf{h}_{u^{\prime}}^{l}$ (11)
+To generate the embedding for node $u$ at layer $l+1$ we perform the following transformation(Hamilton et al., 2017):
+[…]
 where $\mathbf{h}_{u}^{(l)}$ is the node embedding in layer $l$. $\mathbf{W}^{l}_{1}$ and $\mathbf{W}^{l}_{2}$ are trainable weight matrices at layer $l$.
@@
 Building upon this, we compute the score of each partition $p\in\mathcal{P}^{t}$ with respect to the node $v$ selected in Phase 1 using all the nodes in $p$.
-In contrast to predicting a fixed-size score vector corresponding to number of partitions (nazi2019gap; tsitsulin2023graph), the proposed design choice makes the model inductive to the number of partitions.
+In contrast to predicting a fixed-size score vector corresponding to number of partitions (Nazi et al., 2019; Tsitsulin et al., 2023b), the proposed design choice makes the model inductive to the number of partitions.
 Specifically, the number of partitions are not directly tied to the output dimensions of the neural model.
 Having obtained the transformed node embeddings through a Gnn in Eq. 11, we now compute the (unnormalized) score for node $v$ selected in phase 1 for each partition $p\in\mathcal{P}^{t}$ as follows:
-$\operatorname{PartScore}$ $(p,v){=}\operatorname{AGG}(\{\operatorname{MLP}(\sigma(h_{v}|h_{u}))$ (12)
+(12) $\operatorname{PartScore}$ $(p,v){=}\operatorname{AGG}(\{\operatorname{MLP}(\sigma(h_{v}|h_{u}))$
 $\quad\quad\quad\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=p\})$
@@
 Mathematically, the probability of taking action $a^{t}{=}p$ at time step $t$ at state $S^{t}$ is defined as:
-$\pi((a^{t}{=}p)/S^{t})=\frac{\exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$ (13)
-During the course of trajectory of length $T$, we sample action $a^{t}\in\mathbb{R}$ i.e., the assignment of the partition for the node selected in phase 1 at step $t$ using policy $\pi$.
+(13) $\pi((a^{t}{=}p)/S^{t})=\frac{\exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$
+During the course of trajectory of length $T$, we sample action $a^{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathcal{P}}$ i.e., the assignment of the partition for the node selected in phase 1 at step $t$ using policy $\pi$.
 State Transition.
@@
 Specifically, if node $v$ belonged to $i^{th}$ partition at time $t$ and its partition has been changed to $j$ in phase 2, then we apply the below operations in order.
-$P^{t+1}_{i}\leftarrow P^{t}_{i}\setminus v\text{ and }P^{t+1}_{j}\leftarrow P^{t}_{j}\cup v$ (14)
-3.3 Reward, Training & Inference
+(14) $\hskip 43.36243ptP^{t+1}_{i}\leftarrow P^{t}_{i}\setminus\{v\}\text{ and }P^{t+1}_{j}\leftarrow P^{t}_{j}\cup\{v\}\vspace{-0.2in}$
+3.3. Reward, Training & Inference
 Reward.
 Our aim is to improve the value of the overall partitioning objective.
-One way is to define the reward $R^{t}$ at step $t\geq 0$ as the change in objective value of the partitioning i.e $Obj(\mathcal{G},\mathcal{P}^{t})$ at step $t$, i.e., $R^{t}=(Obj(\mathcal{G},\mathcal{P}^{t+1})-Obj(\mathcal{G},\mathcal{P}^{t}))/(Obj(\mathcal{G},\mathcal{P}^{t+1})+Obj(\mathcal{G},\mathcal{P}^{t}))$.
-Here, the denominator term ensures that the model receives a significant reward for even slight enhancements when dealing with a small objective value.
-However, this definition of reward focuses on short-term improvements instead of long-term.
-Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term (sutton2018reinforcement).
+One way is to define the reward $R^{t}$ at step $t\geq 0$ as the change in objective value of the partitioning i.e $Obj(\mathcal{G},\mathcal{P}^{t})$ at step $t$.
+Specifically,
+(15) $R^{t}=\frac{(Obj(\mathcal{G},\mathcal{P}^{t})-Obj(\mathcal{G},\mathcal{P}^{t+1}))}{(Obj(\mathcal{G},\mathcal{P}^{t})+Obj(\mathcal{G},\mathcal{P}^{t+1}))}\cdot\lambda$
+Here $\lambda$ is a hyperparameter that is used to scale the reward.
+The above reward expression incentivizes significant improvements in the objective function.
+This is achieved mathematically by considering both the change and the current value of the objective function.
+This design steers the model towards prioritizing substantial improvements, especially in low objective function regions, ultimately guiding it towards the overall minimum.
+However, above definition of reward focuses on short-term improvements instead of long-term.
+Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term (Sutton and Barto, 2018).
 The discounted rewards are computed as the sum of the rewards over a horizon of actions with varying degrees of importance (short-term and long-term).
 Mathematically,
-$D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}$ (15)
+(16) $D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}$
 where $T$ is the length of the horizon and $\gamma\in(0,1]$ is a discounting factor (hyper-parameter) describing how much we favor immediate rewards over the long-term future rewards.
 The above reward mechanism provides flexibility to our framework to be versatile to objectives of different nature, that may or may not be differentiable.
-This is an advantage over existing neural methods (nazi2019gap; tsitsulin2023graph) where having a differentiable form of the partitioning objective is a pre-requisite.
+This is an advantage over existing neural methods (Nazi et al., 2019; Tsitsulin et al., 2023b) where having a differentiable form of the partitioning objective is a pre-requisite.
 Policy Loss Computation and Parameter Update.
 Our objective is to learn parameters of our policy network in such a way that actions that lead to an overall improvement of the partitioning objective are favored more over others.
-Towards this, we use REINFORCE gradient estimator (williams1992simple) to optimize the parameters of our policy network.
+Towards this, we use REINFORCE gradient estimator (Williams, 1992) to optimize the parameters of our policy network.
 Specifically, we wish to maximize the reward obtained for the horizon of length $T$ with discounted rewards $D^{t}$.
 Towards this end, we define a reward function $J(\pi_{\theta})$ as:
-$J(\pi_{\theta})=\mathbb{E}\big{[}\sum_{t=0}^{T}\left(D^{t}\right)\big{]}$ (16)
-We, then, optimize $J(\pi_{\theta})$ as follows:
-${\nabla J(\pi_{\theta})=\left[\sum_{t=0}^{T}\left(D^{t}\right)\nabla_{\theta}log\pi_{\theta}({a}^{t}/S^{t})\right]}$ (17)
+(17) $J(\pi_{\theta})=\mathbb{E}\big{[}\sum_{t=0}^{T}\left(D^{t}\right)\big{]}$
+We, then, optimize $J(\pi_{\theta})$ via policy gradient (Sutton and Barto, 2018) as follows:
+(18) ${\nabla J(\pi_{\theta})=\left[\sum_{t=0}^{T}\left(D^{t}\right)\nabla_{\theta}log\pi_{\theta}({a}^{t}/S^{t})\right]}$
 Training and Inference.
@@
 During inference, we compute the initial node embeddings, obtain initial partitioning and then run the forward pass of our policy to improve the partitioning objective over time.
-3.4 Time complexity
+3.4. Time Complexity
 The time complexity of NeuroCUT during inference is
-$\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{E}|+k)\times T^{\prime}\big{)}$.
+$\mathcal{O}\big{(}(\alpha\times\beta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times|\mathcal{E}|})+(|\mathcal{E}|+k)\times T^{\prime}\big{)}$.
 Here $\alpha$ is the number of anchor nodes, $\beta$ is number of random walk iterations, $k$ is the number of partitions and $T^{\prime}$ is the number of iterations during inference.
 Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$.
-Further, for sparse graphs $\mathcal{E}=\mathcal{O}(|\mathcal{V}|)$.
+Further, for sparse graphs $|\mathcal{E}|=\mathcal{O}(|\mathcal{V}|)$.
 Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.
-For detailed derivation please see appendix  A.4.
-4 Experiments
+For detailed derivation please see Appendix  A.2.
+4. Experiments
 In this section, we demonstrate the efficacy of NeuroCUT against state-of-the-art methods and establish that:
   • Efficacy and Robustness: NeuroCUT produces the best results over diverse partitioning objective functions. This establishes the robustness of NeuroCUT.
-  • Inductivity: As one of the major strengths, unlike existing neural models such as GAP (nazi2019gap), DMon (tsitsulin2023graph) and MinCutPool (bianchi2020spectral), our proposed method NeuroCUT is inductive on the number of partitions and consequently can generalize to unseen number of partitions.
-4.1 Experimental Setup
-Datasets:
-We use four real datasets for our experiments. They are described below and their statistics are present in the appendix (please see Table 7).
-  • Cora and Citeseer (sen2008collective): These are citation networks where nodes correspond to individual papers and edges represent citations between papers.
+  • Inductivity: As one of the major strengths, unlike existing neural models such as GAP (Nazi et al., 2019), DMon (Tsitsulin et al., 2023b) and MinCutPool (Bianchi et al., 2020), our method NeuroCUT is inductive on the number of partitions and consequently can generalize to unseen number of partitions.
+Code: Our code base is accessible at https://github.com/idea-iitd/NeuroCut.
+4.1. Experimental Setup
+4.1.1. Datasets:
+We use four real datasets for our experiments. They are described below and their statistics are present in Table 2. For all datasets we use their largest connected component.
+  • Cora and Citeseer (Sen et al., 2008): These are citation networks where nodes correspond to individual papers and edges represent citations between papers.
 The node features are extracted using a bag-of-words approach applied to paper abstracts.
-  • Harbin (li2019learning): This is a road network extracted from Harbin city, China.
+  • Harbin (Li et al., 2019): This is a road network extracted from Harbin city, China.
 The nodes correspond to road intersections and node features represent latitude and longitude of a road intersection.
-  • Actor (platonov2023critical): This dataset is based upon Wikipedia data where each node in the graph corresponds to an actor, and the edge between two nodes denotes co-occurrence on the same Wikipedia page.
+  • Actor (Platonov et al., 2023): This dataset is based upon Wikipedia data where each node in the graph corresponds to an actor, and the edge between two nodes denotes co-occurrence on the same Wikipedia page.
 Node features correspond to keywords in Wikipedia pages.
-Partitioning objectives:
+This is a heterophilous dataset (Zheng et al., 2022) where nodes tend to connect to other nodes that are dissimilar.
+  • Facebook (McAuley and Leskovec, 2012): In social network analysis, a $k$-way cut can delineate communities of users with minimal inter-community connections.
+This partitioning unveils hidden social circles and provides insights into how different groups interact within the larger network.
+The Facebook dataset, used for community detection, does not include node features.
+In this dataset, nodes represent users, and edges represent friendships.
+  • Stochastic Block Models (SBM) (Abbe, 2017): SBMs are synthetic networks and produce graphs with communities, where subsets of nodes are characterized by specific edge densities within the community.
+This allows us to validate our method by comparing the identified partitions with the ground truth communities.
+This dataset does not include node features.
+    Table 2. Dataset Statistics
+    Dataset $|\mathcal{V}|$ $|\mathcal{E}|$Features Average Degree Clustering Coefficient Degree Assortativity
+    Cora 2485 5069 1433 4.07 0.23 -0.07
+    CiteSeer 2120 3679 3703 3.47 0.169 0.0075
+    Harbin 6598 10492 2 3.18 0.036 0.22
+    Actor 6198 14879 931 4.82 0.05 -0.048
+    Facebook 1034 26749 - 51.7 0.526 0.431
+    SBM 500 5150 - 20.6 0.18 -0.0059
+4.1.2. Partitioning objectives:
 We evaluate our method on a diverse set of four partitioning objectives described in Section  2, namely Normalized Cut, Balanced Cut, k-MinCut, and Sparsest Cut.
 In addition to evaluating on diverse objectives, we also choose a diverse number of partitions for evaluation, specifically, $k=2,\;5\;\text{and}\;10$.
-Baselines:
+4.1.3. Baselines:
 We compare our proposed method with both neural as well as non-neural methods.
-Neural baselines: We compare with DMon (tsitsulin2023graph), GAP (nazi2019gap), MinCutPool  (bianchi2020spectral) and Ortho (tsitsulin2023graph).
+Neural baselines: We compare with DMon (Tsitsulin et al., 2023b), GAP (Nazi et al., 2019), MinCutPool  (Bianchi et al., 2020) and Ortho (Tsitsulin et al., 2023b).
 DMon is the state-of-the-art neural attributed-graph clustering method.
@@
 Ortho is the orthogonality regularizer described in DMon and MinCutPool.
-Non-neural baselines: Following the settings of DMon (tsitsulin2023graph), we compare with K-means clustering applied on raw node features.
-Other settings:
+We also compare with DRL (Gatti et al., 2022) which solves for Normalized cut at $k=2$ in Appendix A.5
+Non-neural baselines: Following the settings of DMon (Tsitsulin et al., 2023b), we compare with K-means clustering applied on raw node features.
+We also compare with standard graph clustering methods hMetis (Karypis et al., 1997) and Spectral clustering (Ng et al., 2001) in App Sec.
+A.4.
+4.1.4. Other settings:
 We run all our experiments on an Ubuntu 20.04 system running on Intel Xeon 6248 processor with 96 cores and 1 NVIDIA A100 GPU with 40GB memory for our experiments.
-For NeuroCUT we used GraphSage(hamilton2017inductive) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$.
+For NeuroCUT we used GraphSage(Hamilton et al., 2017) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$.
 We used Adam optimizer for training the parameters of our policy network $\pi_{\theta}$.
@@
 The default number of anchor nodes for computing positional embeddings is set to $35$.
+We set $\beta=100$ and $c=0.85$ for Eq.
+6.
+We set scaling factor $\lambda=100$ in eq.
+15.
+Table 3. Results on Normalized Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$.
 Dataset Method $k=2$ $k=5$ $k=10$
-Cora Kmeans 0.65 3.26 7.44
+Cora K-means 0.65 3.26 7.44
 MinCutPool 0.12 0.61 1.65
@@
 NeuroCUT 0.02 0.33 0.92
-CiteSeer Kmeans 0.30 2.35 5.21
+CiteSeer K-means 0.30 2.35 5.21
 MinCutPool 0.10 0.38 1.04
@@
 NeuroCUT 0.02 0.20 0.44
-Harbin Kmeans 0.56 2.34 5.40
+Harbin K-means 0.56 2.34 5.40
 MinCutPool - - -
@@
 NeuroCUT 0.01 0.07 0.28
-Actor Kmeans 0.99 4.00 8.98
+Actor K-means 0.99 4.00 8.98
 MinCutPool 0.55 1.97 4.73
@@
 NeuroCUT 0.17 0.99 4.66
-Table 2: Results on Normalized Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$.
+Table 4. Results on Sparsest Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
 Dataset Method $k=2$ $k=5$ $k=10$
-Cora Kmeans 3.41 14.90 32.52
+Cora K-means 3.41 14.90 32.52
 MinCutPool 0.52 2.44 6.68
@@
 NeuroCUT 0.13 1.46 3.03
-CiteSeer Kmeans 1.53 13.70 22.20
+CiteSeer K-means 1.53 13.70 22.20
 MinCutPool 0.31 1.32 3.62
@@
 NeuroCUT 0.11 0.49 1.19
-Harbin Kmeans 1.91 7.44 17.03
+Harbin K-means 1.91 7.44 17.03
 MinCutPool - - -
@@
 NeuroCUT 0.06 0.23 0.82
-Actor Kmeans 5.43 19.40 40.34
+Actor K-means 5.43 19.40 40.34
 MinCutPool 2.44 8.51 19.81
@@
 NeuroCUT 0.65 2.04 2.88
-Table 3: Results on Sparsest Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
-4.2 Results on Transductive Setting
+4.2. Results on Transductive Setting
 In the transductive setting, we compare our proposed method NeuroCUT against the mentioned baselines, where the neural models are trained and tested with the same number of partitions.
-Tables 2-5 present the results for different partitioning objectives across all datasets and methods.
+Tables 3-6 present the results for different partitioning objectives across all datasets and methods.
 For the objectives under consideration, a smaller value depicts better performance.
@@
 Further, we would also like to point out that in many cases, existing baselines produce “nan” values which are represented by “-” in the result tables.
+This is due to the reason that every partition is not assigned atleast one node which leads to the situation where denominator becomes 00 in normalized, balanced and sparsest cut objectives as defined in Sec.
+2.
 Our method NeuroCUT incorporates dependency during inference and this leads to robust performance.
-Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in methods such as GAP (nazi2019gap).
+Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in methods such as GAP (Nazi et al., 2019).
 Further, unlike other methods, NeuroCUT also incorporates positional information in the form of Lipschitz embedding to better contextualize global node positional information.
 We also observe that the non-neural method K-means fails to perform on all objective functions since its objective is not aligned with the main objectives under consideration.
+Table 5.
+Results on k-MinCut.
+In most of the cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$.
+Values less than $10^{-2}$ are approximated to 0.
 Dataset Method $k=2$ $k=5$ $k=10$
-Cora Kmeans 0.32 0.34 0.61
+Cora K-means 0.32 0.34 0.61
 MinCutPool 0.06 0.12 0.17
@@
 NeuroCUT $\sim$ 0.0 $\sim$ 0.0 0.06
-CiteSeer Kmeans 0.12 0.27 0.43
+CiteSeer K-means 0.12 0.27 0.43
 MinCutPool 0.04 0.08 0.10
@@
 NeuroCUT 0.01 0.02 0.03
-Harbin Kmeans 0.28 0.46 0.54
+Harbin K-means 0.28 0.46 0.54
 MinCutPool 0.0 0.0 0.0
@@
 NeuroCUT $\sim$ 0.0 0.01 0.03
-Actor Kmeans 0.48 0.54 0.80
+Actor K-means 0.48 0.54 0.80
 MinCutPool 0.25 0.35 0.42
@@
 NeuroCUT $\sim$ 0.0 $\sim$ 0.0 $\sim$ 0.0
-Table 4: Results on k-MinCut.
-In most of the cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$.
-Values less than $10^{-2}$ are approximated to 0.
+Table 6. Results on Balanced Cut. In most cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
 Dataset Method $k=2$ $k=5$ $k=10$
-Cora Kmeans 0.68 3.90 7.44
+Cora K-means 0.68 3.90 7.44
 MinCutPool 0.13 0.60 1.62
@@
 NeuroCUT 0.45 0.64 1.08
-CiteSeer Kmeans 0.42 2.64 5.30
+CiteSeer K-means 0.42 2.64 5.30
 MinCutPool 0.09 0.38 1.04
@@
 NeuroCUT 0.07 0.24 0.60
-Harbin Kmeans 0.56 2.37 5.41
+Harbin K-means 0.56 2.37 5.41
 MinCutPool - - -
@@
 NeuroCUT 0.21 0.11 0.27
-Actor Kmeans 1.00 4.06 9.08
+Actor K-means 1.00 4.06 9.08
 MinCutPool 0.55 1.96 4.71
@@
 NeuroCUT 0.59 1.69 4.42
-Table 5: Results on Balanced Cut. In most cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
-4.3 Results on Inductivity to Partition Count
-Dataset Method Normalized Sparsest k-MinCut Balanced
-Cora Kmeans 7.44 32.52 0.61 7.53
-NeuroCUT-T 0.92 3.03 0.06 1.08
-NeuroCUT-I 1.18 5.04 0.02 2.03
-CiteSeer Kmeans 5.21 22.20 0.43 5.30
-NeuroCUT-T 0.44 1.19 0.03 0.60
-NeuroCUT-I 0.63 1.62 0.03 0.62
-Harbin Kmeans 5.40 17.03 0.54 5.41
-NeuroCUT-T 0.28 0.82 0.03 0.27
-NeuroCUT-I 0.27 0.87 0.03 0.29
-Table 6: Inductivity to unseen partition count.
+4.3. Results on Inductivity to Partition Count
+Table 7.
+Inductivity to unseen partition count.
 Here we set the target number of partitions $k=10$.
@@
 In this table, the transductive version of NeuroCUT trained on $k=10$ serves as a point of reference when assessing the quality of the inductive version of NeuroCUT.
-[…]
-[…]
-Figure 2: Results on the initial warm start and the final cut values obtained by NeuroCUT at $k{=}10$.
+Dataset Method Normalized Sparsest k-MinCut Balanced
+Cora K-means 7.44 32.52 0.61 7.53
+NeuroCUT-T 0.92 3.03 0.06 1.08
+NeuroCUT-I 1.18 5.04 0.02 2.03
+CiteSeer K-means 5.21 22.20 0.43 5.30
+NeuroCUT-T 0.44 1.19 0.03 0.60
+NeuroCUT-I 0.63 1.62 0.03 0.62
+Harbin K-means 5.40 17.03 0.54 5.41
+NeuroCUT-T 0.28 0.82 0.03 0.27
+NeuroCUT-I 0.27 0.87 0.03 0.29
+Figure 2.
+Results on the initial warm start and the final cut values obtained by NeuroCUT at $k{=}10$.
 It shows that our neural model NeuroCUT (Final) performs more accurate node and partition selection to optimize the objective function.
 Subsequently, there is a significant difference between the initial and final cut values.
-As detailed in Section 3.2, the decoupling of parameter size and the number of partitions allows NeuroCUT to generalize effectively to an unseen number of partitions.
+As detailed in Section 3.2.2, the decoupling of parameter size and the number of partitions allows NeuroCUT to generalize effectively to an unseen number of partitions.
 In this section, we empirically analyze the generalization performance of NeuroCUT to an unknown number of partitions.
@@
 Note that while the non-neural method such as K-means needs to be re-run for the unseen partitions, NeuroCUT only needs to perform forward pass to produce the results on an unseen partition size.
-4.4 Ablation Studies
+4.4. Ablation Studies
 Initial Warm Start vs Final Cut Values.
@@
 We note that there is a significant difference between the initial and final cut values, which essentially shows the effectiveness of NeuroCUT.
-Figure 3: Node Selection in Phase 1 with Heuristic vs Random: Relative % improvement (gain) in cut values obtained by NeuroCUT when using node selection heuristic against random node selection for $k=5$.
+Figure 3.
+Node Selection in Phase 1 with Heuristic vs Random: Relative % improvement (gain) in cut values obtained by NeuroCUT when using different node selection strategies for $k=5$.
 In most cases, our heuristic finds significantly better cuts than a random node selection procedure.
 […]
-5 Conclusion
+Impact of Clustering Initialization in Warm-start Phase: To understand the importance of different initializations in the warm-up phase(§ 3.1), in Appendix  A.6 we perform an ablation study using three different initialization schemes, namely K-means(default), density-based clustering DBSCAN (Ester et al., 1996), and Random initialization.
+4.5. Impact of Cluster Strength on Performance
+In Section A.8 in appendix, we analyze the impact of cluster strength in networks on the performance of different methods.
+5. Conclusion
 In this work, we study the problem of graph partitioning with node features.
@@
 Notably, our method shows significant performance gains when compared to the state-of-the-art techniques, proving its competence in both inductive and transductive settings.
-In summary, NeuroCUT offers a versatile, inductive, and empirically validated solution for graph partitioning.
+Limitations and Future Directions: Achieving a sub-quadratic computational complexity with an inductive neural method for attributed graph partitioning is an open challenge.
+In NeuroCUT, node selection and perturbation are performed in a sequential fashion.
+One direction to improve the efficiency of NeuroCUT could be batch processing the selection and perturbation of multiple independent nodes simultaneously.
+Another interesting direction could be designing explanations of the neural methods—such as NeuroCUT—for graph combinatorial problems (Kakkad et al., 2023).
+6. Acknowledgement
+We thank Kartik Sharma for the helpful suggestions.
+Rishi Shah acknowledges that the funding for the conference was provided by the CMU GSA Conference Funding.
+Sahil Manchanda acknowledges financial support by GP Goyal grant of IIT Delhi and Qualcomm Innovation Fellowship India.
 Appendix A Appendix
-A.1 Additional Related Work
-In recent years there has been a growing interest in using machine learning techniques to design heuristics for NP-Hard combinatorial problems on graphs.
-This line of work is motivated from the observation that traditional algorithmic approaches are limited to using the same heuristic irrespective of the underlying data distribution.
-Neural approaches, instead, focus on learning the heuristic as a function of the data without deriving explicit algorithms (duong2023deep; jung2022learning; nazi2019gap).
-Significant advances have been made in learning to solve a variety of combinatorial optimization problems such as Set Cover, Travelling Salesman Problem, Influence Maximization etc.
-(khalil2017learning; gcomb; kool2018attention; joshi2019efficient).
-The neural architectural designs in these methods are generally suitable for a specific set of CO problems (mazyavkina2021reinforcement) of similar nature.
-For instance, AM  (kool2018attention) and NeuroLKH (xin2021neurolkh) are targeted towards routing problems.
-GCOMB (gcomb) and LeNSE (ireland2022lense) aim to solve problems that are submodular in nature.
-Owing to these assumptions, the architectural designs in these works are not suitable for our problem of $k$-way inductive graph partitioning with versatile partitioning objectives.
-A.2 Clustering for Initialization
+A.1. Clustering for Initialization
 As discussed in sec 3.1 in main paper, we first cluster the nodes of the graph into $k$ clusters where $k$ is the number of partitions.
-Towards this, we apply K-means algorithm (macqueen1967some) on the nodes of the graph where a node is represented by its raw node features and positional representation i.e $\mathbf{emb_{init}}(u)\;\forall u\in\mathcal{V}$ based upon eq.
+Towards this, we apply K-means algorithm (MacQueen et al., 1967) on the nodes of the graph where a node is represented by its raw node features and positional representation i.e $\mathbf{emb_{init}}(u)\;\forall u\in\mathcal{V}$ based upon eq.
 8.
 Further, we used $L_{\inf}$ norm as the distance metric for clustering.
-A.3 Dataset Statistics
-Table  7 describes the statistics of datasets used in our work. For all datasets we use their largest connected component.
-Dataset $|\mathcal{V}|$ $|\mathcal{E}|$ Number of features
-Cora 2485 5069 1433
-CiteSeer 2120 3679 3703
-Harbin 6598 10492 2
-Actor 6198 14879 931
-Table 7: Dataset Statistics
-A.4 Time complexity of NeuroCUT
-We derive the time complexity of forward pass of NeuroCUT as discussed in sec. 3.4 in main paper.
+A.2. Time Complexity Analysis of NeuroCUT
 (1) First the positional embeddings for all nodes in the graph are computed.
@@
 6).
-This takes $\mathcal{O}(\alpha\times\beta)$ time.
+This takes $\mathcal{O}(\alpha\times\beta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times|\mathcal{E}|})$ time.
 (2) Next, GNN is called to compute embeddings of node.
@@
 Steps 2-4 are repeated for $T^{\prime}$ iterations.
-Hence overall running time is $\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{V}|\times d+k)\times T^{\prime}\big{)}$.
+Hence overall running time is $\mathcal{O}\big{(}(\alpha\times\beta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times|\mathcal{E}|})+(|\mathcal{V}|\times d+k)\times T^{\prime}\big{)}$.
 Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$.
@@
 Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.
+A.3. Time Complexity Comparison
+Table 8 presents the complexities of NeuroCUT and other prominent neural and non-neural baselines algorithms.
+In the table, $|\mathcal{V}|$, $|\mathcal{E}|$, $k$, $d$ are the number of nodes, edges, partitions, and average node degree respectively.
+While some neural algorithms exhibit faster complexity, they require separate training for each partition size ($k$).
+In contrast, NeuroCUT, once trained, can generalize to any value.
+Consequently, for practical workloads, NeuroCUT presents a more scalable option in terms of computation overhead and storage (one model versus separate models for each).
+Table 8. Time complexity of different methods.
+Method Complexity Comments
+DMon ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(d^{2}{\times}|\mathcal{V}|+|\mathcal{E}|)}$ Per iteration complexity
+MinCutPool ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(d^{2}{\times}|\mathcal{V}|+|\mathcal{E}|)}$ Per iteration complexity
+Ortho ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(|\mathcal{V}|{\times}k^{2})}$ Per iteration complexity
+GAP ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(|\mathcal{V}|{\times}k^{2}+|\mathcal{V}|^{2})}$ Per iteration complexity
+Spectral ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(|\mathcal{V}|^{3})}$ Inference complexity
+NeuroCUT ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}o(|\mathcal{V}|^{2})}$ Inference complexity
+The quadratic time complexity of NeuroCUT may pose challenges for very large graphs.
+However, this complexity remains faster than spectral clustering, a widely used graph partitioning algorithm.
+Also, for the baseline neural methods(DMon, MinCutPool, Ortho and GAP), the time complexity provided is per iteration.
+The number of iterations often ranges between 1000 to 2000 and there is no clear understanding of how it varies as a function of the graph (like density, diameter, etc.)
+A.4. Comparison: Non-neural Baselines
+We compare the performance of our method against graph based clustering algorithms hMETIS and Spectral Clustering and partition algorithm Gomory-Hu Tree (Gomory and Hu, 1961).
+Since these methods are not compatible with datasets having raw node features, we compare with two datasets namely Facebook (Leskovec and Mcauley, 2012) and Stochastic Block Model(SBM) (Abbe, 2017) which don’t have node features for this comparision.
+Table 9 compares the performance of NeuroCUT against hMETIS and Spectral Clustering.
+The details of these datasets are presented in Table 2.
+In addition to non-neural methods, we also show the performance of neural methods MinCutPool, DMon, Ortho and GAP on these datasets.
+The performance of NeuroCUT is better than non-neural methods hMETIS, Spectral and Gomory-Hu Tree in most of the cases on Facebook dataset.
+Further, on SBM dataset, it matches the performance of hMETIS and Spectral clustering.
+The SBM dataset has a clear community structure, making it an easy instance for partitioning algorithms, resulting in similar performance across different methods.
+We would also like to highlight that in the case of $k$--$mincut$, Gomory-Hu Tree algorithm generated trivial partitions for these datasets where almost all nodes were assigned the same partition.
+The other neural methods such as MinCutPool, DMon, Ortho, and GAP fail to produce a valid solution in most of the cases.
+This is possibly because these baselines are not robust to perform on datasets without raw node features.
+Overall, NeuroCUT outperforms the baselines on diverse objectives and datasets.
+Table 9. Performance of different methods on Facebook and SBM dataset. Lower values are better. ${}^{\prime}-^{\prime}$ denotes nan.
+Dataset Method Metrics
+Normalized Sparsest Balanced $k$-MinCut
+Facebook NeuroCUT 0.257 4.121 0.972 0.015
+hMETIS 1.15 52.319 1.115 0.200
+Spectral 1.67 4.72 2.459 0.003
+Gomory-Hu Tree 4.00 5.00 4.792 $\sim$0
+MinCutPool - - - 0.023
+DMon - - - $\sim$0
+Ortho - - - 0.05
+GAP - - - 0.025
+SBM NeuroCUT 0.191 3.939 0.191 0.038
+hMETIS 0.191 3.939 0.191 0.038
+Spectral 0.191 3.939 0.191 0.038
+Gomory-Hu Tree 4.003 48.75 4.787 0.0075
+MinCutPool - - - 0.0
+DMon - - - 0.0
+Ortho - - - $\sim$0
+GAP - - - 0.035
+A.5. Comparison with Gatti et al. (Gatti et al., 2022)
+In Table 10, we compare the performance of our method against DRL method proposed by Gatti et al.
+(Gatti et al., 2022) which solves the normalized cut problem for the case where the number of partitions is exactly two.
+We observe that NeuroCUT outperforms DRL on all datasets.
+Table 10. Gatti et al. on Normalized Cut at $k=2$.
+$\dfrac{\text{\bf Method}\rightarrow}{\text{\bf Dataset}\downarrow}$ Cora CiteSeer Harbin Actor SBM Facebook
+NeuroCUT 0.02 0.02 0.01 0.17 0.191 0.257
+Gatti et al. 0.355 0.29 0.13 1.00 0.72 0.92
+A.6. Impact of Clustering Initialization
+To understand the importance of different initializations in the warm-up phase, we perform an ablation study on $2$ datasets namely Cora and CiteSeer using 3 different initialization namely K-means, density-based clustering DBSCAN (Ester et al., 1996), and Random initialization.
+In Table 11 we observe the performance on normalized cut at $k=5$.
+We observe that K-means performs the best in this experiment.
+The improvement observed when using K-means or DBSCAN over Random shows that a good initialization i.e.
+warm-up does help in improving quality of partitions.
+In DBSCAN we set $eps=0.9$ and $min\_samples=2$.
+It is worth noting that DBSCAN’s performance can vary based on the parameter selection.
+Nonetheless, the primary goal of this experiment is to demonstrate the advantageous role of a good initialization, specifically in contrast to random initialization.
+Table 11. Impact of Clustering initialization in Warm-up Phase of NeuroCUT in the normalized cut objective at $k=5$.
+Dataset K-means Random DBSCAN
+Cora 0.33 0.79 0.55
+CiteSeer 0.20 0.51 0.70
+Table 12. Statistics for Normalized Cut (3 runs) for $k=5$
+Dataset Our Heuristic Random Selection Random Selection % Gain
+Avg Best (from best)
+Cora 0.33 $\pm$ 0.0024 0.461 $\pm$ 0.0164 0.445 35
+Citeseer 0.20 $\pm$ 0.00961 0.276 $\pm$ 0.0177 0.258 29
+Harbin 0.073 $\pm$ 0.0027 0.097 $\pm$ 0.0131 0.086 18
+Table 13. Statistics for Balanced Cut (across 3 runs) for $k=5$
+Dataset Our Heuristic Random Selection Random Selection % Gain
+Avg Best (from best)
+Cora 0.642 $\pm$ 0.0065 0.753 $\pm$ 0.0574 0.672 2
+Citeseer 0.248 $\pm$ 0.0049 0.307 $\pm$ 0.039 0.252 0.5
+Harbin 0.1087 $\pm$ 0.0010 0.1313 $\pm$ 0.011 0.1293 19
+Table 14. Number of Nodes Re-assigned after Initial Assignment by NeuroCUT on the Balanced Cut objective.
+Dataset Nodes Perturbed Total Nodes Percentage
+Cora 791 2495 31.7
+Citeseer 413 2120 19.4
+Harbin 187 6598 2.8
+A.7. Impact of Node Selection Procedures
+In Fig.
+3 in main paper, random selection performs well only in Balanced Cuts.
+The objective in Balanced Cuts (Eq.
+3.) is not only a function of the combined weight of cut edges, but also balancing the number of nodes across partitions.
+This additional node-balancing term is not present in the objective functions of the other cut definitions.
+Due to this reduced importance of optimizing the cut value, random does well, since even when an incorrect node is selected, it can still be utilized to keep the partition sizes balanced.
+Moreover, we find that when the warm-up assignment is less accurate and the percentage of nodes re-assigned by NeuroCUT to a different partition is high, a random selection of nodes is closer in efficacy to our heuristic selection.
+In contrast, in Harbin, the percentage of nodes perturbed to a new partition is significantly smaller.
+In this scenario, the probability that random selection will select this small subset is lower and hence amplifying the effectiveness gap between random and heuristic selection.
+A.8. Impact of Cluster Strength on Performance
+To understand the impact of community structure on performance of NeuroCUT, we generate Stochastic Model Block(SBM) graphs with different intra cluster strength.
+In Table 15, we observe the performance of different methods on the normalized cut objective at $k=5$.
+The baseline methods hMETIS and Spectral Clustering perform worse when the strength within communities is low indicated by Intra Cluster Edge Probability and Clustering Coefficient values.
+Although NeuroCUT outperforms or matches existing methods in all cases, however, the gap between baselines and NeuroCUT increases significantly when the community structure is not strong.
+Further, the neural baselines DMon, MinCutPool and Ortho generated ‘nan’ values for this experiment.
+Table 15. Performance on SBM Dataset
+SBM Dataset Statistics Normalized Cut
+Intra Cluster Edge Probability Clustering Coefficient $|\mathcal{V}|$ $|\mathcal{E}|$ NeuroCUT hMETIS Spectral Clustering
+1 0.005 0.003 495 555 0.3089 0.331 0.379
+2 0.01 0.0029 525 661 0.3434 0.5889 0.4259
+3 0.2 0.184 500 5150 0.1912 0.1912 0.1912
+4 0.4 0.3822 500 10102 0.1153 0.1153 0.1153
+A.9. Impact of $\beta$ on NeuroCUT
+We study the impact of $\beta$ parameter(eq.
+6) which is the number of iterations in random walk with restart.
+In Table 16 we present the performance of NeuroCUT using different $\beta$ on normalized-cut objective at $k=5$.
+We observe that NeuroCUT improves with more iterations as $\beta$ increases and then its performance stabilizes.
+Table 16. Impact of $\beta$ parameter on NeuroCUT on normalized cut at $k=5$ on the Cora dataset.
+$\beta$ Normalized Cut
+1 2.12
+3 1.57
+10 0.33
+50 0.33
+100 0.33
+A.10. Stability Across Multiple Runs
+Table 17 shows the Mean and Standard deviation of NeuroCUT for Normalized cut at $k=5$ for 3 runs. The Std Dev. is lower than the Mean showing the stability of NeuroCUT across multiple runs.
+Table 17. Mean and Standard Deviation for different datasets on Normalized Cut at $k=5$ for NeuroCUT
+Dataset Mean $\pm$ Std Dev
+Cora $0.337\pm 0.0024$
+Citeseer $0.20\pm 0.00961$
+Harbin $0.073\pm 0.0027$
+Actor $0.97\pm 0.0389$